### v1

# Bit Threads: From Entanglement to Geometric Entropies

Abstract
In this work, we attempt to construct bit threads configurations for various backgrounds using expressions from CPS formalism. We find that when the Ryu-Takayanagi surface is same as horizon, such expressions are sufficient. In other cases, it differs by gradient of a harmonic function. We explore its relation to Wald and differential entropy, and re-express the first law of entanglement entropy in terms of bit threads. Inclusion of quantum effects imposes some constraints on the bulk entanglement via the dominant energy condition. We also apply our method to ascertain bit thread configuration in certain dynamical spacetime.
1 Introduction
Holographic entanglement entropy has emerged as a useful concept in gauge/gravity duality, providing a crucial bridge between quantum information theory and the geometry of spacetime [1]. It takes us one step closer to understanding various issues regarding the quantum nature of gravity, such as the black hole information problem and many aspects of AdS/CFT correspondence. The Ryu-Takayanagi (RT) formula, equates the entanglement entropy $S_{A}$ of a boundary region $A$ in the dual conformal field theory (CFT) to the area of a minimal codimension-2 surface $m(A)$ anchored on $\partial A$ within a constant time slice of the bulk asymptotically Anti-de Sitter (AdS) spacetime [1, 2]:
$S_{A}=\dfrac{Area(m(A))}{4G_{N}}.$ (1.1)
Recently, Freedman and Headrick [3] introduced a novel, flow-based method for computing holographic entanglement entropy (HEE) that bypasses the need for prior identification of the RT surface. In this approach, the entanglement entropy is obtained by maximising the flux of a divergenceless vector field over the boundary region. These vector field lines, termed “bit threads”, can be interpreted as representing the quantum correlations (or “Bell pairs”) that connect different parts of the boundary [3]. Bit threads, thus provide a geometric representation of entanglement, offering an alternative perspective to the traditional minimal surface prescription. In this framework, the concept of Bell pairs plays a central role in interpreting the physical meaning of bit threads. Bell pairs, which represent maximally entangled two-level quantum systems, serve as the fundamental units of entanglement. In the bit-thread picture, each thread can be thought of as carrying the information of one Bell pair. In other words, the number of bit threads that connect a given boundary region to its complement reflects the number of Bell pairs shared between these regions, thereby quantifying the entanglement entropy. This interpretation is motivated by the idea that entanglement entropy is a measure of the nonlocal quantum correlations between subsystems. By associating each bit thread with a Bell pair, the flow-based prescription provides an intuitive geometrical representation of these correlations. The maximisation of bit thread flux, which is equivalent to finding the minimal surface in the traditional Ryu-Takayanagi prescription, effectively counts the number of Bell pairs across the boundary. While the identification of each individual thread with a specific Bell pair is a heuristic picture, it offers an alternate way to visualise the abstract notion of entanglement in holographic theories. This perspective not only reinforces the connection between geometry and quantum information but also opens avenues for exploring how modifications to the entanglement structure might reflect in the underlying spacetime geometry. In this flow prescription, the HEE is written as,
$S_{A}=\max_{\upsilon}\int_{A}\upsilon,$ (1.2a)
subject to the constraints:
$\nabla_{a}\upsilon^{a}=0\,;\quad|\upsilon|\leq\dfrac{1}{4G_{N}}.$ (1.2b)
Here, $\upsilon^{a}$ is the bit threads vector field and at the RT surface $\upsilon^{a}|_{RT}=\mathrm{C}n^{a}$, where $n^{a}$ is the unit normal of the RT surface and $\mathrm{C}=1/4G_{N}$. The equivalence between this max-flow picture and the minimal-area RT formula is guaranteed by a continuous version of the max-flow min-cut theorem, where the RT surface $m(A)$ emerges as the bottleneck, or minimal cut, whose capacity dictates the maximum possible flux (see Appendix A for a technical review).
This elegant flow-based perspective has turned out to be a useful prescription for a variety of problems. It has been generalised to covariant settings, introduced by Headrick and Hubeny [4], accommodates the dynamics of entanglement in non-static spacetimes by employing a family of divergenceless flows consistent with the extremal surface prescription in the Hubeny-Rangamani-Takayanagi (HRT) framework [5] and beyond that. Additionally, bit threads have been applied to study various information theoretic quantities in holography. They have provided insights into the monogamy of mutual information [6], a fundamental constraint on the sharing of entanglement among multiple parties, as well as the entanglement of purification [7, 8, 9], which quantifies correlations in mixed quantum states. The Riemannian bit threads have also been extended to cases like higher curvature gravity [10], quantum bit threads [11, 12], PEE threads [13], and entanglement threads [14] to name some of the few.
Major thrust of this paper is to explore the relationship between bit threads and the rigorous framework of the covariant phase space (CPS) formalism [15, 16]. Building on prior work that constructed perturbative bit threads using CPS methods [17], we investigate whether the HEE bit thread field can be directly identified with conserved currents derived from CPS. Bit threads, constructed in this way has been referred to as canonical bit threads. This formalism provides a covariant phase space approach, and its integration with bit threads offers a promising avenue for deeper insights into the relationship between geometric and entropic quantities in holography. But there are certain challenges when we directly want to obtain the bit thread vector field from some form derived using the CPS formalism. We find out that while CPS inspired vector fields $v^{a}_{cps}$ are divergenceless, these generally do not satisfy the norm bound and saturation conditions required for bit threads. An extra term $v_{\mathrm{d}}^{a}$ (which does not change the cohomology class of the dual form) needs to be incorporated in many cases, except for certain situations like Rindler-AdS geometries where the RT surface coincides with the horizon. But once we obtain the vector field in Rindler like coordinates, we can use coordinate transformations to find a valid thread configuration in the other coordinate system where we initially needed to add $v^{a}_{\mathrm{d}}$ (like Poincaré) without any prior knowledge of the RT surface.
This connection inspires us to explore whether the flow paradigm can account for other geometric entropy concepts. We argue that the Wald entropy (associated with black hole and cosmological horizons) and differential entropy (associated with spacetimes with a hole) can indeed be naturally expressed as the flux of specific CPS-derived flows linked to the relevant Killing vectors. Furthermore, the CPS method allows us to reformulate the first laws governing variations of these entropies in terms of associated thermodynamic flows, offering a unified geometric perspective for entropy.
Finally, we address the incorporation of quantum effect, discussing the constraint it provides on the bulk entanglement via energy conditions. We also outline the challenges and potential strategies for constructing covariant bit threads in certain dynamical spacetimes, useful for studying processes like thermalisation and black hole evaporation.
The paper is organised as follows. We discuss the construction of a flow from the CPS formalism in section 2. Then we move towards a non-holographic direction to reinterpret the black hole entropy, entropy of cosmological horizons and differential entropy in terms of these threads in section 3. In the section 4, we discuss the first laws of the all the above mentioned entropies in terms of bit threads. The quantum corrections and their relation to the dominant energy condition are discussed in section 5. Section 6 considers dynamical settings. Finally, we conclude with the summary of the work and possible future directions in section 7.
2 Bit threads and CPS
In general, it is difficult to obtain a bit thread configuration unless the given geometry is very symmetric. There have been some proposals of constructing these vector fields using geodesics [18], unit normals to the RT surfaces, level setting approach [17], or as magnetic flows using a modified Biot-Savart law [19, 20]. Usually, we need to know the metric to do calculations. This motivates us to seek a more fundamental approach, potentially one less reliant on the explicit metric and more connected to the geometry, intrinsic symmetries and conserved structures of the gravitational theory. The CPS formalism, which operates naturally with differential forms and conserved currents, presents a promising avenue [^1] Another approach is to use calibrations, in which the RT surface is realised as a special Lagrangian cycle [21].. So, a natural question arises: Can we construct flows in a more canonical way using the CPS method? [^2] An attempt to construct bit thread configuration using the Noether current form was made in [22]. The authors of [17] have already explored this direction for perturbative bit threads, terming them ‘canonical bit threads’. Here, we investigate how this idea can be extended to non-perturbative cases as well. The main trick is to use parametric variations instead of general perturbations within the CPS framework, which works equally well. Sometimes this is referred to as the ‘solution phase space’ method [23]. For completeness, a short review of the CPS formalism is provided in appendix B. Although there are various versions of the CPS formalism, we will mainly refer to that espoused by Iyer-Wald [15], as all of them lead to the same entropy formula, which is of our primary interest. First, we will sketch a general idea of how to get the bit thread vector field from the CPS, and then we will look into its applicability.
In CPS formalism, the spacetime is considered to be a $d$-dimensional manifold and the phase space is a symplectic manifold. The Hamiltonian then is given by integrating the symplectic current $\boldsymbol{\omega}$ (See (B.3)-(B.8)) over a Cauchy slice. This construction is not only restricted to the Hamiltonian, but is true for any conserved charge associated with a specific set of transformations generated by the symmetry generators. It can be seen if one proceeds as follows. The symplectic current is a closed (d1)(d-1)-form, i.e.,
$\mathrm{d}\boldsymbol{\omega}(\delta_{1}\Phi,\delta_{2}\Phi,\Phi)=0,$ (2.1)
for any two perturbations satisfying the linearised equations of motion. Then, on the phase space $\cal{P}$}, one can locally define a (d2;1)(d-2;1)-form $\mathbf{k}_{\zeta}$ for any transformation generator $\zeta$ as
$\boldsymbol{\omega}(\delta\Phi,\delta_{\zeta}\Phi,\Phi)=\mathrm{d}\mathbf{k}_{\zeta}(\delta\Phi,\Phi).$ (2.2)
This equation is called the ‘fundamental identity’ of the CPS formalism. The variation of a conserved charge associated with $\zeta$ is given by,
$\delta H_{\zeta}(\Phi)=\int_{\Sigma}\mathrm{d}\mathbf{k}_{\zeta}(\delta\Phi,\Phi)=\int_{\partial\Sigma}\mathbf{k}_{\zeta}(\delta\Phi,\Phi).$ (2.3)
This is basically the equation (B.8) in terms of the (d2)(d-2)-form $\mathbf{k}$.
2.1 Codim-2 Current
As mentioned earlier, there are different versions of the CPS formalisms, we will look into the codimension-2 current $\mathbf{k}_{\zeta}$ [^3] Sometimes it is called the ‘surface charge’, but we will stick to the codim-2 current as it is convenient to interpret when we construct a flow using it. for two of them namely the Iyer-Wald formalism and the covariant phase space with boundaries.
Codim-2 Current in Iyer-Wald CPS
In the Iyer-Wald formalism [15], the symplectic current is given by,
$\boldsymbol{\omega}_{\rm{IW}}(\delta_{1}\Phi,\delta_{2}\Phi,\Phi)=\delta_{1}\mathbf{\Theta}_{\rm{IW}}(\delta_{2}\Phi,\Phi)-\delta_{2}\mathbf{\Theta}_{\rm{IW}}(\delta_{1}\Phi,\Phi),$
and for a diffeomorphism invariant theory, it becomes
$\boldsymbol{\omega}_{\rm{IW}}(\delta\Phi,\mathcal{L}_{\xi}\Phi,\Phi)=\delta\mathbf{\Theta}_{\rm{IW}}(\mathcal{L}_{\xi}\Phi,\Phi)-\mathcal{L}_{\xi}\mathbf{\Theta}_{\rm{IW}}(\delta\Phi,\Phi),$ (2.4)
with $\xi$ being the diffeomorphism generator. Here $\zeta$ is replaced with $\xi$ to differentiate between a diffeomorphism generator and any general transformation. To obtain an explicit expression for $\mathbf{k}$, the other ingredient is the Noether current, which is defined in terms of the diffeomorphism generator $\xi$, symplectic potential $\mathbf{\Theta}_{\rm{IW}}$ and the Lagrangian $\boldsymbol{\mathrm{L}}$ as given in equation (B.6): $\boldsymbol{\mathrm{J}}_{\xi}\equiv\mathbf{\Theta}_{\rm{IW}}(\mathcal{L}_{\xi}\Phi,\Phi)-\xi.\boldsymbol{\mathrm{L}}$. Then the variation of the Noether current $\boldsymbol{\mathrm{J}}_{\xi}$ yields
$\delta\boldsymbol{\mathrm{J}}_{\xi}=\delta\mathbf{\Theta}_{\rm{IW}}(\mathcal{L}_{\xi}\Phi,\Phi)-\mathcal{L}_{\xi}\mathbf{\Theta}_{\rm{IW}}(\delta\Phi,\Phi)+\mathrm{d}\left(\xi.\mathbf{\Theta}_{\rm{IW}}(\delta\Phi,\Phi)\right),$
where Cartan’s first structure equation has been used. Substituting this in the expression of $\boldsymbol{\omega}_{\rm{IW}}$, one gets
$\boldsymbol{\omega}_{\rm{IW}}$ $=\delta\boldsymbol{\mathrm{J}}_{\xi}-\mathrm{d}(\xi.\mathbf{\Theta}_{\rm{IW}})=\delta(\mathrm{d}\boldsymbol{\mathrm{Q}}_{\xi})-\mathrm{d}(\xi.\mathbf{\Theta}_{\rm{IW}}),$
$=\mathrm{d}(\delta\boldsymbol{\mathrm{Q}}_{\xi}-\xi.\mathbf{\Theta}_{\rm{IW}}).$ (2.5)
Here, $\boldsymbol{\mathrm{Q}}_{\xi}$ is the Noether charge defined in (B.7). Now, comparing the above equation with the fundamental identity (2.2), the codimension-2 current for diffeomorphism invariant theories is given by,
$\mathbf{k}_{\xi}=\delta\boldsymbol{\mathrm{Q}}_{\xi}-\xi.\mathbf{\Theta}_{\rm{IW}}.$ (2.6)
If one can find a (d1)(d-1)-form $\boldsymbol{B}$ such that $\mathbf{\Theta}_{\rm{IW}}=\delta\boldsymbol{B}$ then the above equation can be rewritten as
$\mathbf{k}_{\xi}=\delta\boldsymbol{\mathit{j}}_{\xi},$ (2.7)
where the codim-2 form $\boldsymbol{\mathit{j}}_{\xi}$ is given by,
$\boldsymbol{\mathit{j}}_{\xi}=\boldsymbol{\mathrm{Q}}_{\xi}-\xi.\boldsymbol{B}\;.$ (2.8)
This current form is our main ingredient to construct a configuration for bit threads.
Codim-2 current in CPSB
In the covariant phase space formalism with boundaries [24], one starts with an action of the theory which has an extra boundary Lagrangian (B.9) and follows the similar procedure as stated earlier. In this case, the (pre)-symplectic current is defined as,
$\boldsymbol{\omega}_{\rm{HW}}=\delta\left(\mathbf{\Theta}_{\rm{HW}}-\mathrm{d}\boldsymbol{\mathrm{C}}\right)|_{\tilde{\mathcal{P}}}$ (2.9)
where, $\tilde{\mathcal{P}}$ denotes the pre-phase space. The symplectic current is closed: $\mathrm{d}\boldsymbol{\omega}_{\rm{HW}}=0$. Pre-symplectic form is defined as the symplectic current integrated over a Cauchy surface, $\tilde{\mathbf{\Omega}}_{\rm{HW}}=\int_{\Sigma}\boldsymbol{\omega}_{\rm{HW}}$. Then, the change in Hamiltonian is given by
$\delta H_{\xi}=-X_{\xi}.\tilde{\Omega}_{HW}=\int_{\Sigma}\left(-X_{\xi}.\boldsymbol{\omega}_{\rm{HW}}\right),$ (2.10)
where, $\xi^{\mu}$ is the diffeomorphism generator and $X_{\xi}$ is a vector field defined on the configuration space given by, $X_{\xi}=\int d^{d}x\mathcal{L}_{\xi}\Phi^{i}(x)\frac{\delta}{\delta\Phi^{i}}$. In CPSB, the Noether current is expressed as
$\boldsymbol{\mathrm{J}}_{\xi}\equiv X_{\xi}.\mathbf{\Theta}_{\rm{HW}}-\xi.\boldsymbol{\mathrm{L}},$
and the Noether charge is $\boldsymbol{\mathrm{Q}}_{\xi}$ which satisfies $\boldsymbol{\mathrm{J}}_{\xi}=\mathrm{d}\boldsymbol{\mathrm{Q}}_{\xi}$ as $\mathrm{d}\boldsymbol{\mathrm{J}}_{\xi}=0$. Substituting these in the r.h.s of (2.10), one gets
$X_{\xi}.\boldsymbol{\omega}_{\rm{HW}}=\delta\boldsymbol{\mathrm{J}}_{\xi}+\mathrm{d}\left(\delta_{\xi}\boldsymbol{\mathrm{C}}-\delta\left[X_{\xi}.\boldsymbol{\mathrm{C}}\right]-\xi.\mathbf{\Theta}_{\rm{HW}}\right).$
Now, operating exterior derivative on the above quantity gives us
$\mathrm{d}\left(-X_{\xi}.\boldsymbol{\omega}_{\rm{HW}}\right)=\delta\mathrm{d}\boldsymbol{\mathrm{J}}_{\xi}+\mathrm{d}^{2}\left(\delta_{\xi}\boldsymbol{\mathrm{C}}-\delta\left[X_{\xi}.\boldsymbol{\mathrm{C}}\right]-\xi.\mathbf{\Theta}_{\rm{HW}}\right)=0,$
as $\boldsymbol{\mathrm{J}}_{\xi}$ is locally an exact form and $\mathrm{d}^{2}=0$ by definition. It follows immediately that $X_{\xi}.\boldsymbol{\omega}$’ is a closed form. Locally it can be thought as an exact form and can be expressed as an exterior derivative of some codimension-2 quantity, i.e., $X_{\xi}.\boldsymbol{\omega}_{\rm{HW}}=\mathrm{d}\tilde{\mathbf{k}}_{\xi}$, where $\tilde{\mathbf{k}}_{\xi}$ is given by,
$\tilde{\mathbf{k}}_{\xi}=\delta\boldsymbol{\mathrm{Q}}_{\xi}+\delta_{\xi}\boldsymbol{\mathrm{C}}-\delta\left[X_{\xi}.\boldsymbol{\mathrm{C}}\right]-\xi.\mathbf{\Theta}_{\rm{HW}}\,.$
We can further simplify it by using the boundary conditions and Cartan’s equation, which results in
$\tilde{\mathbf{k}}_{\xi}=\delta\left(\boldsymbol{\mathrm{Q}}_{\xi}+\xi.\boldsymbol{{\ell}}-X_{\xi}.\boldsymbol{\mathrm{C}}\right).$ (2.11)
As seen from the above equation, we can find a current $\tilde{\boldsymbol{\mathit{j}}_{\xi}}$ such that $\tilde{\mathbf{k}}_{\xi}=\delta\tilde{\boldsymbol{\mathit{j}}_{\xi}}$, which is expressed as
$\tilde{\boldsymbol{\mathit{j}}_{\xi}}=\left(\boldsymbol{\mathrm{Q}}_{\xi}+\xi.\boldsymbol{{\ell}}-X_{\xi}.\boldsymbol{\mathrm{C}}\right).$ (2.12)
This codimension-2 current can be used to construct various conserved charges. Now we will look how these current forms can be used to construct a flow for bit threads.
2.2 Bit threads from CPS
In the previous subsection we saw that how one can find the (d2)(d-2)-form $\mathbf{k}_{\xi}$ (and $\boldsymbol{\mathit{j}}_{\xi}$) from the covariant phase space formalism. Here, we will show how to construct a flow from this form that is a potential bit threads vector field. Both of the above mentioned CPS formalisms are equally valid for this construction as they lead to the same entropy. We will mainly follow the IW prescription. The codimension2-2 current $\mathbf{k}_{\xi}$ is given by
$\mathbf{k}_{\xi}=\delta\boldsymbol{\mathrm{Q}}_{\xi}-\xi.\mathbf{\Theta}_{\rm{IW}}\,$,
for any diffeomorphism generator $\xi^{\mu}$. From hereafter we will be omitting the suffix ‘IW’ to simplify the notations. In the covariant phase space formalism [15, 23] these codim-2 currents are used to obtain variations of different conserved charges in a black hole background which lead to the renowned first law of black hole thermodynamics (see appendix B.3 for a short review). We can extend this method to obtain the conserved charges by using the (d2)(d-2)-form $\boldsymbol{\mathit{j}}_{\xi}$ (if present) instead of the usual $\mathbf{k}_{\xi}$. We mainly focus on the exact symmetries (like isometry) of the theory. When $\xi^{\mu}$ is a Killing vector (or any exact symmetry generator), the symplectic current $\boldsymbol{\omega}(\delta\Phi,\mathcal{L}_{\xi}\Phi,\Phi)$ vanishes, i.e.,
$\boldsymbol{\omega}(\delta\Phi,\mathcal{L}_{\eta}\Phi,\Phi)=0=\mathrm{d}\mathbf{k}_{\eta}(\delta\Phi,\Phi)=\mathrm{d}\delta\boldsymbol{\mathit{j}}_{\eta}(\delta\Phi,\Phi),$
where we have denoted $\eta$’ as the set of all Killing vectors $\xi$’s. The above equation implies that $\mathbf{k}_{\eta}(\delta\Phi,\Phi)$ is a closed form and so is $\boldsymbol{\mathit{j}}_{\eta}$. Then in the presence of a Riemannian metric $g_{ab}$ ($a,b=1,\cdots,d-1$), we can have a divergenceless vector field $v^{a}$ from both the closed forms $\mathbf{k}_{\eta}(\delta\Phi,\Phi)$ and $\boldsymbol{\mathit{j}}_{\eta}$. We will use the latter one to obtain a vector field using the map:
$v_{cps}^{a}=g^{ab}(\star\boldsymbol{\mathit{j}}_{\eta})_{b},$ (2.13)
where, $\star$’ is the Hodge dual defined on a Cauchy surface. Again, we can write [25],
$\mathrm{d}\boldsymbol{\mathit{j}}_{\eta}=(\nabla_{a}v_{cps}^{a})\boldsymbol{\epsilon},$ (2.14)
where $\boldsymbol{\epsilon}$ is the volume form of the Cauchy slice. As $\boldsymbol{\mathit{j}}_{\eta}$ is a closed form, it implies that the vector field is indeed divergenceless (i.e., $\nabla_{a}v_{cps}^{a}=0$) everywhere.444 On the uniqueness of $\boldsymbol{v_{cps}^{a}}$ : The fundamental identity (2.2) allows us to define the codim-2 current $\mathbf{k}_{\xi}$ from the symplectic current, but this definition involves some ambiguities that we need to address. The symplectic potential $\mathbf{\Theta}$ itself is not uniquely defined [15]. Adding boundary terms to the Lagrangian changes $\mathbf{\Theta}$ but leaves the symplectic current $\boldsymbol{\omega}=\delta\mathbf{\Theta}$ unchanged. More generally, we can modify $\mathbf{\Theta}\rightarrow\mathbf{\Theta}-\mathrm{d}\mathbf{B}$, which transforms the symplectic current as $\boldsymbol{\omega}\rightarrow\boldsymbol{\omega}+\mathrm{d}\boldsymbol{\omega}_{B}$. These are called “residual ambiguities” and reflect our freedom in choosing boundary terms [26]. However, for exact symmetries (such as Killing symmetries), these ambiguities cancel out. When we transform the symplectic current $\boldsymbol{\omega}\rightarrow\boldsymbol{\omega}+\mathrm{d}\boldsymbol{\omega}_{B}$, the codim-2 current transforms as: $\mathbf{k}_{\xi}\rightarrow\mathbf{k}_{\xi}+\boldsymbol{\omega}_{B}(\delta\Phi,\mathcal{L}_{\xi}\Phi,\Phi)$ For Killing vectors or generators of exact symmetries, the Lie derivative $\mathcal{L}_{\xi}\Phi$ equals the field variation $\delta_{\xi}\Phi$, so the additional term vanishes and $\mathbf{k}_{\xi}$ remains unchanged. This means the vector field $v_{cps}^{a}$ obtained through the CPS construction is uniquely defined (up to the usual degrees of freedom in $\mathbf{k}_{\xi}$) when we work with exact symmetries. Equivalently we can have the relation at the boundary of a Cauchy surface as,
$\boldsymbol{\mathit{j}}_{\eta}\Big{|}_{\partial\Sigma}=(n_{a}v_{cps}^{a})\tilde{\boldsymbol{\epsilon}}$, (2.15)
where, $\tilde{\boldsymbol{\epsilon}}$ is volume form on its boundary. We use the (d2)(d-2)-form $\mathbf{k}$ to construct a vector field which will help us to study thermodynamics of a gravitational system. We will discuss this briefly in the section 4.
A central requirement of the bit thread formalism is that $v^{a}$ must be divergenceless. The vector fields derived from the CPS formalism using the codim-2 form $\mathbf{k}_{\xi}$ (or $\boldsymbol{\mathit{j}}_{\xi}$) automatically fulfil this condition. This suggests that imposing the remaining bit thread constraints onto these CPS-derived flows might provide a canonical construction for bit threads. To have a better understanding of the connection between these two formalisms, we must re-examine the properties of the codim-2 form $\mathbf{k}_{\xi}$. Although $\mathbf{k}_{\xi}$ resolves ambiguities present in the symplectic current (as explained in the footnote 4), it can be expressed upto an exact form. This freedom in $\mathbf{k}_{\xi}$ can be utilised to satisfy the properties of the bit threads. So we can write the current form as
$\mathbf{k}_{\xi}\rightarrow\mathbf{k}_{\xi}+\mathrm{d}\mathbf{k}_{B_{\xi}}.$
This tells us that $\mathbf{k}_{\xi}$ belongs to the cohomology class $\left[\mathbf{k}\right]=\{\mathbf{k}_{\xi}\,|\,\mathrm{d}\mathbf{k}_{\xi}=0,\,\mathbf{k}_{\xi}\sim\mathbf{k}_{\xi}+\mathrm{d}\mathbf{k}_{B_{\xi}}\}\in H_{dR}^{d-2}(\mathcal{M})$. Similarly, the current form $\boldsymbol{\mathit{j}}_{\xi}$ belongs to the cohomology class $\left[\boldsymbol{\mathit{j}}\right]\in H_{dR}^{d-2}(\mathcal{M})$ given by,
$\boldsymbol{\mathit{j}}_{\xi}\sim\boldsymbol{\mathit{j}}_{\xi}^{\prime}\quad s.t.\quad\boldsymbol{\mathit{j}}_{\xi}^{\prime}-\boldsymbol{\mathit{j}}_{\xi}=\mathrm{d}\boldsymbol{\mathit{j}}_{B_{\xi}}\,.$ (2.16)
The same ensues to the vector field obtained from the above current form:
$v^{a}=g^{ab}(\star\boldsymbol{\mathit{j}}_{\xi})_{b}+g^{ab}(\star\mathrm{d}\boldsymbol{\mathit{j}}_{B_{\xi}})_{b}=v^{a}_{cps}+v_{\mathrm{d}}^{a},$ (2.17)
where, $v^{a}_{\mathrm{d}}$ is the vector field corresponding to the co-exact $1$-form $\star\mathrm{d}\boldsymbol{\mathit{j}}_{B_{\xi}}$ and is divergence free. Now, if the given manifold (or the domain of our interest) is simply connected, $v^{a}_{\mathrm{d}}$ can be expressed as a gradient of some function $\varphi(x)$. So, the term $v_{\mathrm{d}}^{a}=\nabla^{a}\varphi$ is like a $U(1)$ gauge freedom [^5] We thank Mohd Ali for emphasising the importance of this term and for valuable discussions.. This freedom helps us to satisfy the boundary conditions for bit threads in some cases. We are mainly interested in the regions of the spacetime manifold where the domain is simply connected. Then, we can say that $v^{a}$ and $v^{a}_{cps}$ are from the same equivalence class :
$v^{a}]=\Bigl{\{}v^{a}_{cps}\,|\,v^{a}_{cps}\sim v^{a}_{cps}+\nabla^{a}\varphi\,\Bigl{\}}\,\quad\;~\text{with}\;\nabla^{2}\varphi=0\,.$ (2.18)
Evidently, $v^{a}$ is a divergenceless vector field and we will show that this becomes a valid bit thread configuration once we impose the conditions (1.2b). If the manifold has non-trivial topologies (when the manifold is not simply connected), the bit threads field $v^{a}$ and the vector derived from CPS $(v^{a}_{cps})$ differ by a divergenceless vector field $v^{a}_{\mathrm{d}}$. Therefore, in general the equivalence class of $v^{a}_{cps}$ can be written as: $v^{a}_{gen}]=\{v^{a}_{cps}\,|\,\,v^{a}_{cps}\sim v^{a}_{cps}+v^{a}_{\mathrm{d}}\,\}$ with $\nabla_{a}v^{a}_{\mathrm{d}}=0$.
The bit threads vector field that we obtained using the (d2)(d-2)-form $\boldsymbol{\mathit{j}}$ arising from the CPS formalism is called ‘canonical bit thread’. Although this formalism does not require any notion of metric until we switch to the notion of a vector: the map (2.13) precisely use the metric of a Cauchy surface to construct the desired $v_{cps}^{a}$ from the differential form $\boldsymbol{\mathit{j}}$. And it is essential to invoke the non-trivial $v^{a}_{\mathrm{d}}$ to satisfy all the conditions of bit threads. For examples we will look into AdS3 spacetime in two different coordinate systems.
Poincaré AdS3
Here we will discuss the construction of bit threads in Poincaré $\rm{AdS}_{3}$ background for some boundary interval and show how naively applying the $v_{cps}^{a}$ fails to reproduce the HEE and we do need to add a non-trivial vector field $v_{\mathrm{d}}^{a}$ to satisfy all the conditions of bit threads . The metric in Poincaré coordinates is given by,
$ds^{2}=\dfrac{L^{2}}{z^{2}}\left(-dt^{2}+dz^{2}+dx^{2}\right),$ (2.19)
where, $L$ is the AdS radius, $x$ is the boundary coordinate and the boundary CFT is at $z=0$. We will use the bulk modular flow of the $t=0$ slice: $\xi^{bulk}_{modular}|_{t=0}=\left(\dfrac{\pi}{R}[R^{2}-z^{2}-x^{2}]\partial_{t},0,0\right)$, which is a Killing vector of the spacetime. To find $\mathbf{k}_{\xi}$, we use the expression [23],
$k^{\mu\nu}_{\xi}=\dfrac{1}{16\pi G_{N}}\left[\left(\xi^{\nu}\nabla^{\mu}h-\xi^{\nu}\nabla_{\sigma}h^{\mu\sigma}+\xi_{\sigma}\nabla^{\nu}h^{\mu\sigma}+\dfrac{1}{2}h\nabla^{\nu}\xi^{\mu}-h^{\delta\nu}\nabla_{\delta}\xi^{\mu}\right)-(\mu\leftrightarrow\nu)\right],$ (2.20)
where, $\xi^{\mu}$ is a Killing vector, $h^{\mu\nu}=g^{\mu\sigma}g^{\nu\gamma}\,\delta g_{\sigma\gamma}$ is the metric perturbation and $h\equiv h^{\mu}_{\mu}=g_{\mu\nu}h^{\mu\nu}$ is its trace. These indices run from 0 to d1d-1 for a $d$ dimensional spacetime. The codimension-2 form is then given by,
$\mathbf{k}_{\xi}=\dfrac{\sqrt{-g}}{(d-2)!\,2!}\,\epsilon_{\mu\nu\sigma_{1}\cdots\sigma_{d-2}}\,k^{\mu\nu}_{\xi}dx^{\sigma_{1}}\wedge\cdots\wedge dx^{\sigma_{d-2}}.$
To obtain the explicit expression of $k^{\mu\nu}_{\xi}$, we need to choose some variation of the metric (2.25). There are different ways to calculate the variation of the metric $\delta g_{\mu\nu}$. One of the most useful and also convenient for our case is to use the parametric variation denoted as
$\hat{\delta}g_{\mu\nu}(P,Q)=\dfrac{\partial g_{\mu\nu}}{\partial P}\hat{\delta}P+\dfrac{\partial g_{\mu\nu}}{\partial Q}\hat{\delta}Q,$
where, $P$ and $Q$ are some parameters of the metric. For the present case (2.19), the only constant we have is the AdS radius $L$. So, the parametric variation looks like $\hat{\delta}g_{\mu\nu}(L)=\dfrac{\partial g_{\mu\nu}}{\partial L}\hat{\delta}L$. Using this parametric variation, we find the only non-zero component of the form $\mathbf{k}_{\xi}$ :
$k^{tz}_{\xi}=-\dfrac{\hat{\delta}L}{8G_{N}}\dfrac{z}{L^{3}}\left(\dfrac{R^{2}-x^{2}}{R}\right).$
We can also find another codim-2 current $\boldsymbol{\mathit{j}}_{\xi}$ such that $\mathbf{k}_{\xi}=-\hat{\delta}\boldsymbol{\mathit{j}}_{\xi}$ given by
$\mathit{j}^{tz}_{\xi}=\dfrac{1}{16G_{N}}\dfrac{z}{L^{2}}\left(\dfrac{R^{2}-x^{2}}{R}\right).$ (2.21)
Now using the relation (2.13), we can find a divergenceless vector field $v_{cps}^{a}$ to be [^6] We thank Manish Ramchander for pointing out a mistake in this expression at the earlier stage of the work.,
$v_{cps}^{a}=\left(v_{cps}^{z},v_{cps}^{x}\right)=\dfrac{\mathfrak{B}}{4G_{N}}\;\dfrac{z^{2}}{2R}\big{(}R^{2}-x^{2},0\big{)},$ (2.22)
where, $\mathfrak{B}$ is an overall constant and is a function of $L$. We have kept the factor $1/4G_{N}$ outside the constant to keep track of the constraints (1.2b). From this expression it is clear that $v_{cps}^{a}$ is divergenceless. But it fails to be proportional to the $n^{a}$ on the RT surface. Recall the RT surface in Poincaré coordinates is given by $x^{2}+z^{2}=R^{2}$, $R$ being the radius of the RT surface. Then the unit normal is $n^{a}=(n^{z},n^{x})=\frac{z_{m}}{R}(z_{m},x_{m})$, where $z_{m}$ and $x_{m}$ are the coordinates on the minimal surface. One can easily check that $v_{cps}^{a}\neq n^{a}/4G_{N}$ on the RT surface. So $v^{a}_{cps}$ alone can not reproduce the HEE. But we can use the gauge freedom $v^{a}_{\mathrm{d}}=\nabla^{a}\varphi$ associated with $v^{a}_{cps}$ to make it a bit thread vector field. To find $v_{\mathrm{d}}^{a}$, we first solve $\nabla_{a}\nabla^{a}\varphi=0$ in the Poincaré background, and then substitute $v^{a}_{\mathrm{d}}$ with the gradient of $\varphi$. Solving of the Laplace’s equation with proper boundary conditions gives us the harmonic function
$\varphi=\dfrac{l^{2}_{\rm{AdS}}}{4G_{N}R}\left(z-\dfrac{\mathfrak{B}z^{3}}{6}+\mathfrak{B}\dfrac{zx^{2}}{2}+\dfrac{x\sqrt{R^{2}-z^{2}}}{z^{2}}\left(z-\mathfrak{B}z^{3}\right)\right).$ (2.23)
Taking gradient of the above function and putting together with (2.22), we find the total vector field,
$v^{a}=\dfrac{1}{4G_{N}}\left(\dfrac{z^{2}}{R}+\dfrac{\mathfrak{B}z^{2}}{2R}(R^{2}-x^{2}-z^{2}),\mathfrak{B}\dfrac{z^{3}x}{R}+\dfrac{\sqrt{R^{2}-z^{2}}}{R}\left(z-\mathfrak{B}z^{3}\right)\right),$ (2.24)
which reduces to $n^{a}$ on the RT surface and also satisfies the property $|v|\leq 1/4G_{N}$. So this is a flow which correctly reproduces the HEE of a single interval in Poincaré AdS. To interpret the different parts of $v^{a}$, we can say that the $v_{cps}^{a}$ store all the information regarding the conserved charges of the theory as it is derived from the symplectic current, and $\nabla^{a}\varphi$ helps to satisfy the constraints of bit threads. From this example, it is clear that $v^{a}$ and $v^{a}_{cps}$ are from the same equivalence class, i.e., $v^{a}\sim v^{a}_{cps}$.
A similar treatment which we used in the Poincaré background can be done for other backgrounds as well. As an example we discuss the case of D11-brane in the appendix C. There can be other ways to find the $v_{\mathrm{d}}^{a}$ in different spacetimes which are more convenient to use than the method presented here. In case of Poincaré $\rm{AdS}$, we can directly obtain a flow solely using the methods of characteristics starting from the divergenceless condition. We briefly discuss it in appendix D.
Rindler AdS3
In the previous subsection, we have seen that it was required to use the freedom presented in the definition of $v_{cps}^{a}$ to satisfy all the conditions to be a flow even in a simple geometry like Poincaré AdS. So, a natural question arises: Can we find geometries where $v_{cps}^{a}$ is sufficient to be a flow without the aid of its gauge freedoms? To answer this question, we need to go to the reference frame of an observer whose causal boundary is bounded by the RT surface. Then in the natural coordinates of that observer, we can obtain a valid bit thread vector field from the CPS formalism. One can always map a spherical region of the boundary CFT to this kind of wedge in the hyperbolic foliation of the dual AdS spacetime [27]. For example, consider an accelerated observer in $\rm{AdS}_{3}$ spacetime. The reference frame of the observer is the Rindler-$\rm{AdS}_{3}$ geometry,
$ds^{2}=-\left(\dfrac{\rho^{2}}{L^{2}}-\mathtt{a}^{2}\right)d\tau^{2}+\left(\dfrac{\rho^{2}}{L^{2}}-\mathtt{a}^{2}\right)^{-1}d\rho^{2}+\rho^{2}du^{2}\,,\qquad u\in\mathbb{R}\,,~\tau\sim\tau+2\pi i,$ (2.25)
where $\mathtt{a}$ is the Rindler acceleration parameter, $\tau$ is the Rindler time and $\rho=\mathtt{a}L$ is the horizon. We use the parametric variation $\hat{\delta}g_{\mu\nu}(\mathtt{a})=\dfrac{\partial g_{\mu\nu}}{\partial\mathtt{a}}\hat{\delta}\mathtt{a}$ to obtain the only non-zero component of $k_{\xi}^{ab}$,
$k_{\xi}^{\tau\rho}=\dfrac{1}{8\pi G_{N}}\dfrac{\mathtt{a}\hat{\delta}\mathtt{a}}{\rho}.$
Here, we have chosen the Killing vector to be $\xi^{\mu}=(\partial_{\tau},0,0)$. We can find a $\boldsymbol{\mathit{j}}$ such that $\mathbf{k}_{\xi}=\hat{\delta}\boldsymbol{\mathit{j}}_{\xi}$, and its non-trivial component is given by,
$j_{\xi}^{\tau\rho}=\dfrac{1}{16\pi G_{N}}\dfrac{\mathtt{a}^{2}}{\rho}.$ (2.26)
This form is defined in the whole spacetime. By projecting it on a Cauchy surface, and using the map (2.13), we find
$v^{\rho}=\dfrac{1}{4G_{N}}\dfrac{\mathtt{a}\,L}{\rho}\sqrt{\dfrac{\rho^{2}}{L^{2}}-\mathtt{a}^{2}}\,.$
By definition, the thread lines start at the boundary subregion (here the whole boundary) and end at the complement of that subregion and horizon or any other codimension-2 surface present in the spacetime. But the above vector $v^{\rho}$ is directed towards the boundary as that is the direction in which $\rho$ increases. So the actual vector field will be
$v^{a}=\left(-v^{\rho},v^{u}\right)=-\dfrac{1}{4G_{N}}\left(\dfrac{\mathtt{a}\,L}{\rho}\sqrt{\dfrac{\rho^{2}}{L^{2}}-\mathtt{a}^{2}},0\right).$ (2.27)
The above vector field is divergenceless by definition and satisfies the necessary conditions for a valid bit thread configuration: $|v|\leq 1/4G_{N},$ and $v^{a}=-n^{a}/4G_{N}$ on the horizon [^7] Here we are having an extra minus sign because the unit normal is pointing outward at the horizon whereas the direction of the bit thread is towards the horizon. of Rindler space (i.e., the RT surface). These conditions ensure consistency with the bit thread formalism (see equation (A.1)), allowing us to conclude that this vector field constitutes a physically admissible bit thread flow.
Rindler AdS to Poincaré AdS
Here we will show how to obtain a bit thread vector field in Poincaré coordinates from the vector (2.27) by using the coordinate transformations presented in appendix E. First, we have to make sure that the metrics presented earlier and in the appendix are same to use the transformations. Comparing the metrics (2.25) and (E.8), we set the acceleration parameter $\mathtt{a}=1$ in (2.25) which yields
$ds^{2}=-\left(\dfrac{\rho^{2}}{L^{2}}-1\right)d\tau^{2}+\left(\dfrac{\rho^{2}}{L^{2}}-1\right)^{-1}d\rho^{2}+\rho^{2}du^{2},$
and the vector field (2.27) becomes
$v^{a}=\left(-v^{\rho},v^{u}\right)=-\dfrac{1}{4G_{N}}\left(\dfrac{L}{\rho}\sqrt{\dfrac{\rho^{2}}{L^{2}}-1},0\right)=\dfrac{1}{4G_{N}}\left(\dfrac{L}{\rho}\sqrt{\dfrac{\tilde{\rho}^{2}}{L}},0\right).$ (2.28)
We can write the above vector in different bases as
$\vec{v}=v^{a}\partial_{a}=\underbrace{\left(-v^{\rho}\partial_{\rho},v^{u}\partial_{u}\right)}_{Rindler\,AdS}=\underbrace{\left(v^{z}\partial_{z},v^{x}\partial_{x}\right)}_{Poincar\acute{e}\,AdS}$
From (2.28), we see that $v^{u}=0$. So, in Poincaré coordinates the vector will be
$v^{a}=(v^{z},v^{x})=-v^{\rho}(z,x)\left(\dfrac{\partial z}{\partial\rho},\dfrac{\partial x}{\partial\rho}\right).$ (2.29)
To find $v^{\rho}(z,x)$, we need to find $\tilde{\rho}$. Using the relations (E.11) and (E.12), we get
$\tilde{\rho}=\dfrac{1}{2z}(L^{2}-z^{2}-x^{2}).$ (2.30)
Substituting it and the expression of $\rho=\rho(z,x)$ as (E.11a) in (2.28) yields
$\bar{v}^{\rho}(z,x)=4G_{N}v^{\rho}(z,x)=\dfrac{\tilde{\rho}}{\rho}=\dfrac{(L^{2}-z^{2}-x^{2})}{\sqrt{\left(L^{2}+x^{2}+z^{2}\right)^{2}-4L^{2}x^{2}}}.$ (2.31)
Now, using the coordinate transformations (E.12), we find
$\dfrac{\partial z}{\partial\rho}=-\dfrac{2z^{2}}{\sqrt{\left(L^{2}+x^{2}+z^{2}\right)^{2}-4L^{2}x^{2}}}\left(\dfrac{L^{2}-x^{2}+z^{2}}{L^{2}-z^{2}-x^{2}}\right),$ (2.32a)
$\dfrac{\partial x}{\partial\rho}=-\dfrac{4z^{2}}{\sqrt{\left(L^{2}+x^{2}+z^{2}\right)^{2}-4L^{2}x^{2}}}\left(\dfrac{z\,x}{L^{2}-z^{2}-x^{2}}\right).$ (2.32b)
Putting the above relations along with (2.31) in (2.29) gives us a vector field in the Poincaré coordinates,
$v^{a}=\dfrac{2}{4G_{N}}\left(\dfrac{z}{\sqrt{\left(L^{2}+x^{2}+z^{2}\right)^{2}-4L^{2}x^{2}}}\right)^{2}\left(L^{2}-x^{2}+z^{2},2z\,x\right).$
After rearranging a little bit, we find
$v^{a}=\dfrac{1}{4G_{N}L}\left(\dfrac{2Lz}{\sqrt{\left(L^{2}+x^{2}+z^{2}\right)^{2}-4L^{2}x^{2}}}\right)^{2}\left(\dfrac{L^{2}-x^{2}+z^{2}}{2L},\dfrac{z\,x}{L}\right)=\dfrac{1}{4G_{N}L}v^{a}_{geodesic}\,$, (2.33)
where, $v^{a}_{geodesic}$ is the symmetric flow constructed using geodesics in [18]. This vector field is divergenceless, obeys the norm bound $|v|\leq\frac{1}{4G_{N}L}$, and at the RT surface reduces to $v^{a}|_{RT}=\frac{n^{a}}{4G_{N}L}$. Evidently, it is a valid thread configuration. It is surprising that only using the coordinate transformations we found a thread configuration in Poincaré background without any prior knowledge of the RT surface, whereas in constructing the symmetric flow from geodesics the knowledge of the RT surface were required as shown in [18].
Figure 1: Flow lines (black curves with arrows) for bit threads obtained from transforming a valid thread configuration in AdS-Rindler to Poincaré. The gray horizontal line denotes the boundary and the red semicircle is the RT surface corresponding to the boundary subregion (shown in blue) $A\in[-2,2]$.
So far, we have found a thread configuration for a fixed length interval $L$, but in general there can be any size of interval in the boundary CFT. In the bulk theory we have corresponding RT surfaces of different radii. These RT surfaces can be related via the isometries of the gravity theory. As discussed in appendix E, one such simple isometry is boost (E.13) in the embedding plane $(Y_{-1},Y_{2})$. We can find the corresponding vector field in this case as well by proceeding the similar way as above. In the boosted case the expression (2.30) becomes
$\tilde{\rho}=\dfrac{e^{\beta}}{2z}\left(e^{-2\beta}L^{2}-z^{2}-x^{2}\right).$ (2.34)
Using this and (E.16a) in (2.28) yields
$\bar{v}^{\rho}(z,x)=\dfrac{\tilde{\rho}}{\rho}=\dfrac{\left(e^{-2\beta}L^{2}-z^{2}-x^{2}\right)}{\sqrt{\left(e^{-2\beta}L^{2}+x^{2}+z^{2}\right)^{2}-4e^{-2\beta}L^{2}x^{2}}}.$ (2.35)
And using the coordinate relations (E.17), we get
$\dfrac{\partial z}{\partial\rho}=-\dfrac{2e^{-\beta}z^{2}}{\sqrt{\left(e^{-2\beta}L^{2}+x^{2}+z^{2}\right)^{2}-4e^{-2\beta}L^{2}x^{2}}}\left(\dfrac{e^{-2\beta}L^{2}-x^{2}+z^{2}}{e^{-2\beta}L^{2}-z^{2}-x^{2}}\right),$ (2.36a)
$\dfrac{\partial x}{\partial\rho}=-\dfrac{4e^{-\beta}z^{2}}{\sqrt{\left(e^{-2\beta}L^{2}+x^{2}+z^{2}\right)^{2}-4e^{-2\beta}L^{2}x^{2}}}\left(\dfrac{z\,x}{e^{-2\beta}L^{2}-z^{2}-x^{2}}\right).$ (2.36b)
Combining all these in (2.29), we find
$v^{a}=\dfrac{L}{4G_{N}}\left(\dfrac{2e^{-\beta}z}{\sqrt{\left(e^{-2\beta}L^{2}+x^{2}+z^{2}\right)^{2}-4e^{-2\beta}L^{2}x^{2}}}\right)^{2}\left(\dfrac{e^{-2\beta}L^{2}-x^{2}+z^{2}}{2e^{-\beta}L},\dfrac{z\,x}{e^{-\beta}L}\right).$
Now, if we identify $R=e^{-\beta}L$ to be the radius of the entangling surface in the boundary theory, we end up with a vector field proportional to $v^{a}_{geodesic}$:
$v^{a}=\dfrac{1}{4G_{N}L}\left(\dfrac{2Rz}{\sqrt{\left(R^{2}+x^{2}+z^{2}\right)^{2}-4R^{2}x^{2}}}\right)^{2}\left(\dfrac{R^{2}-x^{2}+z^{2}}{2R},\dfrac{z\,x}{R}\right).$ (2.37)
The flow lines of this vector field are shown in fig. 1. The HEE corresponding to a boundary subregion $A$ in terms of the above thread configuration can be written as,
$S_{A}=L\int_{m\sim A}v\,.$ (2.38)
3 Extensions to other entropies
Bit threads can be related to other notions of entropy. In this section we specifically attempt to relate it to black hole (BH) entropy and differential entropy. Since the CPS formalism is commonly applied in the context of black hole thermodynamics [16, 15], a natural extension is to derive a flow representation for black hole entropy. The same thing can be done for differential entropy [28, 29], which is defined in both holographic and non-holographic spacetimes.
3.1 Black hole entropy
The entropy of a black hole has been a topic of interest for a long time since Bekenstein [30] to the present day. Among many different approaches, the approach to define the black hole entropy as the Noether charge using the CPS formalism developed by Wald [15, 16] is a well established and standard notion of entropy, known as the ‘Wald entropy’. We will construct a flow corresponding to the Wald entropy of a black hole. Following the procedure described in section 2.2, we can find a divergenceless vector field. The only difference will be the norm bound at the RT surface. As we are interested in the black hole entropy, the norm bound should be on a minimal/extremal surface homologous to the black hole horizon (or, more specifically the bifurcation surface). So the relevant condition becomes $v^{a}=-n_{\mathcal{H}}^{a}/4G_{N}$ at the black hole horizon, $n_{\mathcal{H}}^{a}$ being the unit normal to the horizon. Similar to the AdS-Rindler case, the minus sign emphasises that the direction of $v^{a}$ is towards the horizon [^8] If we work with a Poincaré like coordinate system, then $v^{a}$ is naturally towards the horizon and we do not need to put the minus sign.. Then, the BH entropy is given by,
$S_{BH}=\int_{\cal{B}}v,$ (3.1)
[…]
BTZ black hole
The metric of a BTZ black hole is given by,
$ds_{\rm{BTZ}}^{2}=-\mathcal{N}^{2}dt^{2}+\mathcal{N}^{-2}dr^{2}+r^{2}\left(\mathcal{N}^{\phi}dt+d\phi\right)^{2}$
where, $\mathcal{N}$ denotes the ‘lapse’ function and $\mathcal{N}^{\phi}$ is the angular shift given by,
$\mathcal{N}^{2}=-M+\dfrac{r^{2}}{l^{2}}+\dfrac{J^{2}}{4r^{2}}\;;\quad\mathcal{N}^{\phi}=-\dfrac{J}{2r}.$
Here, ‘$M$’ and ‘$J$’ are the mass and angular momentum of the BTZ black hole respectively; $l$ is the AdS radius. For simplicity, we will consider a non-rotating BTZ black hole, i.e., $J=0$ and $\mathcal{N}^{\phi}=0$. Now, the metric becomes,
$ds_{\rm{BTZ}}^{2}=-\mathcal{N}^{2}dt^{2}+\mathcal{N}^{-2}dr^{2}+r^{2}d\phi^{2}$ (3.2)
with $\mathcal{N}^{2}=-M+\dfrac{r^{2}}{l^{2}}$. The horizon of this black hole is at $r_{\mathcal{H}}=l\sqrt{M}$. Consider the metric to be the function of mass of the BTZ black hole, i.e., $g_{ab}=g_{ab}(M)$. Then the parametric variation becomes,
$\hat{\delta}g_{ab}(M)=\dfrac{\partial g_{ab}}{\partial M}\hat{\delta}M.$
The different components of this variation are given by,
$\hat{\delta}g_{tt}=\hat{\delta}M\;\;;\;\;\hat{\delta}g_{rr}=\mathcal{N}^{-4}\hat{\delta}M=\left(\dfrac{r^{2}}{l^{2}}-M\right)^{-2}\hat{\delta}M;\;\;\hat{\delta}g_{\phi\phi}=0.$
So, the only non-zero components of $h^{ab}$ are,
$h^{tt}=g^{tt}g^{tt}\hat{\delta}g_{tt}=\mathcal{N}^{-4}\hat{\delta}M=\left(\dfrac{r^{2}}{l^{2}}-M\right)^{-2}\hat{\delta}M;$
and,
$h^{rr}=g^{rr}g^{rr}\hat{\delta}g_{rr}=\mathcal{N}^{4}\mathcal{N}^{-4}\hat{\delta}M=\hat{\delta}M.$
Evidently, $h^{ab}$ is traceless, $h=h^{a}_{a}=0$. Now, the antisymmetric nature of $k^{ab}$ tells us that all the diagonal components are zero, i.e., $k^{tt}=k^{rr}=k^{\phi\phi}=0$. The only probable non-zero components are $k^{tr}$, $k^{t\phi}$, $k^{\phi r}$, and their antisymmetric pairs. Using (2.20) it is straightforward to get the non-zero component $k^{tr}=\dfrac{1}{16\pi G_{N}r}\hat{\delta}M$, for the Killing vector $\xi^{a}=(\partial_{t},0,0)$. For this parametric variation, we can easily find an expression for $\boldsymbol{\mathit{j}}_{\xi}$ given by,
$j_{\xi}^{tr}=\dfrac{1}{16\pi G_{N}r}M.$ (3.3)
[…]
$v^{a}=\left(-v^{r},v^{\phi}\right)=-\dfrac{1}{4G_{N}}\left(\dfrac{\mathcal{N}M}{\kappa\,r},0\right),$ (3.4)
where, $\kappa=r_{\mathcal{H}}/l^{2}$ is the surface gravity of the horizon. This vector field correctly reproduces the entropy of the BTZ black hole, indicating that it is a flow corresponding to a valid bit thread configuration for the black hole (or Wald) entropy. Generalisation to the rotating BTZ black hole is straightforward.
Schwarzschild-AdS black hole
A similar procedure can be done for other black holes. Here, we consider a $(d+2)$-dimensional Schwarzschild-AdS black hole. The metric is given by,
$ds^{2}_{\rm{Sch}}$ $=-f(r)dt^{2}+\dfrac{dr^{2}}{f(r)}+r^{2}d\Omega^{2}_{d}$
$=-f(r)dt^{2}+\dfrac{dr^{2}}{f(r)}+r^{2}d\phi^{2}+r^{2}d\Omega^{2}_{d-1},$ (3.5)
where, $f(r)=1-\dfrac{2M}{r}+r^{2}$ with AdS radius $L=1$. Considering spherical symmetries we can suppress the last term in (3.1). Now the metric becomes,
$ds^{2}_{\rm{Sch}}=-f(r)dt^{2}+\dfrac{dr^{2}}{f(r)}+r^{2}d\phi^{2}.$ (3.6)
The rest of the procedure is similar to the last case. Like the previous one, Schwarzschild-AdS is also a solution of the Einstein-Hilbert gravity. Then we can use (2.20) to have the codim-2 current. Parametric variation gives us the different components of the variations of the metric as follows,
$\hat{\delta}g_{tt}=\dfrac{2}{r}\hat{\delta}M~;~\hat{\delta}g_{rr}=\dfrac{2}{rf(r)^{2}}\hat{\delta}M~;~\hat{\delta}g_{\phi\phi}=0.$
These lead us to the metric perturbations,
$h^{tt}=g^{tt}g^{tt}\hat{\delta}g_{tt}=\dfrac{2}{rf(r)^{2}}\hat{\delta}M,\qquad h^{rr}=g^{rr}g^{rr}\hat{\delta}g_{rr}=\dfrac{2}{r}\hat{\delta}M.$
In this case, $h^{ab}$ is traceless: $h^{a}_{a}=h=0$. Now returning to $\mathbf{k}_{\xi}$, the antisymmetric nature tells us that $k^{tt}=k^{rr}=k^{\phi\phi}=0$. Moreover, we find that $k^{r\phi}=0=k^{t\phi}$. The only non-zero component is $k^{tr}=\dfrac{1}{8\pi G_{N}r^{2}}\hat{\delta}M$. And the $\boldsymbol{\mathit{j}}_{\xi}$ is given by,
$j_{\xi}^{tr}=\dfrac{1}{8\pi G_{N}r^{2}}M.$ (3.7)
This $d$-form or the corresponding vector $v^{a}$ related to it via (2.13) gives the black hole entropy upon integration on a codimension-2 surface homologous to the horizon (or more specifically the bifurcation point).
We can think of these flows as the conserved current corresponding to different gravitational charges, and the divergenceless condition is the condition for conservation of some charges similar to the continuity equation.
3.2 Entropy of cosmological horizon
Cosmological horizons, like those in de Sitter space, represent another important class of causal boundaries in spacetime with associated thermodynamic properties. Similar to black hole horizons, cosmological horizons emit Hawking radiation with a characteristic temperature and possess an entropy proportional to their area. Here, we explore how the bit thread formalism can be extended to describe the entropy of cosmological horizons. The de Sitter spacetime in static coordinates is described by the metric:
$ds^{2}=-\left(\mathfrak{a}^{2}-\frac{r^{2}}{L_{dS}^{2}}\right)dt^{2}+\left(\mathfrak{a}^{2}-\frac{r^{2}}{L_{dS}^{2}}\right)^{-1}dr^{2}+r^{2}d\Omega^{2}_{d-2},$ (3.8)
where, $\mathfrak{a}$ is some parameter, $L_{dS}$ is the de Sitter radius related to the cosmological constant by $\Lambda=\frac{(d-1)(d-2)}{2L_{dS}^{2}}$. The cosmological horizon is located at $r=\mathfrak{a}L_{dS}$, and its entropy is given by the Gibbons-Hawking formula:
$S_{dS}=\frac{{\cal{A}}(horizon)}{4G_{N}}=\frac{\Omega_{d-2}L_{dS}^{d-2}}{4G_{N}},$ (3.9)
where $\Omega_{d-2}$ is the volume of a unit (d2)(d-2)-sphere.
To construct a bit thread representation of this entropy, we follow the same approach used for black hole horizons. We use the timelike Killing vector $\xi^{a}=(\partial_{t})^{a}$ and apply the covariant phase space formalism to obtain the codimension-2 form $\mathbf{k}_{\xi}$. For parametric variations of the metric with respect to the de Sitter radius $L_{dS}$, we can find:
$k_{\xi}^{tr}=\frac{1}{8\pi G_{N}}\dfrac{\mathfrak{a}\hat{\delta}\mathfrak{a}}{r},$
which leads to the current form:
$j_{\xi}^{tr}=\frac{1}{16\pi G_{N}}\dfrac{\mathfrak{a}^{2}}{r}.$ (3.10)
Using the mapping in equation (2.13), we can construct a divergenceless vector field:
$v^{a}=\left(-v^{r},v^{\Omega}\right)=-\dfrac{1}{4G_{N}}\left(\frac{\mathfrak{a}L_{dS}}{r}\sqrt{\mathfrak{a}^{2}-\frac{r^{2}}{L_{dS}^{2}}},0\right).$
In the standard dS static patch metric we generally take $\mathfrak{a}=1$ and it can be obtained by rescaling the coordinates in (3.8). Then the above vector field becomes
$v^{a}=-\dfrac{1}{4G_{N}}\left(\frac{L_{dS}}{r}\sqrt{1-\frac{r^{2}}{L_{dS}^{2}}},0\right)$ (3.11)
This vector field satisfies the necessary conditions for a valid bit thread configuration: it is divergenceless, and proportional to the unit normal $n_{\cal{C}}^{a}$ on the cosmological horizon. When integrated over the horizon, it correctly reproduces the Gibbons-Hawking entropy or Wald entropy in more general cases.
An interesting feature of cosmological horizons is that they are observer-dependent. Different observers in de Sitter space perceive different cosmological horizons. This observer dependence is reflected in the bit thread configuration, which is tied to the specific static patch associated with a particular observer.
3.3 Differential entropy
In the AdS/CFT correspondence, entanglement entropy in the boundary theory is computed via the Ryu-Takayanagi prescription, which asserts that the entanglement entropy of a boundary region is equal to the minimum area of a surface in the bulk spacetime. This gives us important insights into how quantum information is encoded in spacetime itself. From the gravitational point of view, entanglement entropy generalises from individual intervals to overlapping boundary intervals. Each of them has an entanglement entropy given by the Ryu-Takayanagi formula. Taking the continuum limit of differences between these entropies, one introduces differential entropy [29]:
$S_{diff}=\int d\lambda\frac{\partial S\big{(}\gamma_{L}(\lambda),\gamma_{R}(\lambda)\big{)}}{\partial\lambda}.$
For particular circumstances, this integral provides the gravitational entropy of a bulk “hole”. Namely, differential entropy happens to be equal to the area of the hole divided by $4G_{N}$,
$S_{diff}=\frac{{\cal{A}}(\rm hole)}{4G_{N}}.$ (3.12)
This discovery indicates a profound relationship between measures of boundary entanglement and bulk geometry. The differential entropy agrees with the Bekenstein-Hawking entropy, which is an implication of the area law, strengthens the holographic principle and demonstrates that the intricate entanglement pattern on the boundary contains the same information as the bulk geometry. Moreover, differential entropy is a versatile tool for the investigation of bulk causal wedges and the reconstruction of bulk regions from boundary data. It provides insightful perspectives on how entanglement warps spacetime and enhances our knowledge of gravitational thermodynamics and holography.
Here, we show that the differential entropy corresponding to a hole in a spacetime can be obtained using the flow prescription. The vector field is derived from CPS formalism with constraints $|v|\leq 1/4G_{N}$ and $|v^{a}n_{a}|=1$ at the boundary of the hole. This $v^{a}$ is divergenceless by definition. So, we can rewrite the differential entropy in terms of the threads as
$S_{diff}=\int_{\mathfrak{b}}v,$ (3.13)
where $\mathfrak{b}$ is the boundary of the hole and $v=\sqrt{\mathfrak{h_{b}}}\,|n_{a}v^{a}|$. We will look into the explicit construction of the vector field in spherical Rindler space.
Spherical Rindler space
The concept of spherical Rindler space generalises the foundational framework of conventional Rindler space that describes a family of observers undergoing radial acceleration away from a shared central point in Minkowski spacetime, establishing causal isolation from a spherical region of radius $R_{0}$ [28]. Unlike linear Rindler acceleration which produces planar horizons, this spherical generalisation creates an observationally inaccessible domain bounded by a perfectly symmetrical horizon surface. The spherical Rindler horizon manifests as a topological boundary separating the causally connected exterior region from the interior zone that remains perpetually hidden to all accelerated observers. This horizon’s geometrical properties maintain exact spherical symmetry, with its characteristic scale being precisely determined by the parameter $R_{0}$ that defines the radius of the inaccessible central region. The $(d+1)$-dimensional metric is given by [28]:
$ds^{2}=-r^{2}dt^{2}+dr^{2}+(R_{0}+r\cosh{t})^{2}d\Omega_{d-1}^{2}\,$, (3.14)
where, $d\Omega_{d-1}^{2}$ denotes the metric on a unit (d1)(d-1)-sphere. This metric structure reveals the intrinsic relationship between the acceleration parameter $R_{0}$, the radial coordinate $r$, and the temporal component that governs the proper time experience of the spherical Rindler observers. The coordinate singularity at $r=R_{0}$ mathematically encodes the causal boundary beyond which no information can propagate to the accelerating observers, establishing this radius both as geometrical and informational horizon. The authors of [28] showed that this horizon gives rise to a notion of entropy proportional to the area of the horizon, which was named as ‘differential entropy’. To obtain the entropy associated with this hole, we generally take the near horizon limit $r\rightarrow 0$. The metric approaches in the limit
$ds^{2}=-r^{2}dt^{2}+dr^{2}+R_{0}^{2}\,d\Omega_{d-1}^{2}\,.$ (3.15)
Now we can find a constant time slice in this geometry and do the Iyer-Wald formalism to find the entropy of the hole which is essentially given by (3.12). We can also find a flow or bit thread configuration by proceeding in the way discussed in section 2.2. In the near horizon geometry we take the radius of the hole $R_{0}$ as the parameter to be used for the parametric variation of the metric. Using (2.20) and (2.7), we find ,
$k^{tr}_{\xi}=\dfrac{\hat{\delta}\mathfrak{c}}{16\pi G_{N}r}\;;\quad j^{tr}_{\xi}=\dfrac{\mathfrak{c}}{8\pi G_{N}r},$ (3.16)
for the near horizon geometry of the spherical Rindler space. Here $\mathfrak{c}$ is a function of $R_{0}$. Now using the relations (2.7) and (2.13), we find the vector field,
$v^{a}=\left(-v^{r},v^{\Omega}\right)=-\dfrac{1}{4G_{N}}\left(\dfrac{\mathfrak{c}}{\kappa},0\right)\,.$ (3.17)
Notice that the vector field is constant along the $r$ direction and the minus sign tells that it ends on the horizon of the hole. This reproduces the entropy of the hole as given by (3.12). So we can see that it is possible to construct a flow based picture for differential entropy. Moreover, from the above construction we can say that the CPS formalism can also be applied to obtain the entropy of hole in a spacetime.
Spherical AdS-Rindler space
Spherical AdS-Rindler space can be viewed as a spherical Rindler space embedded in an asymptotically AdS spacetime. This space is holographically dual to the ultraviolet part of the boundary field theory, described by a finite time strip. [29]. This enables us to study a hole in a spacetime holographically. Analogous to the spherical Rindler space, this spacetime also has a family of accelerated observers whose worldlines are causally disconnected from a spherical hole in AdS. The metric of such spacetime is given by [29],
$ds^{2}=l_{AdS}^{2}\left(-\sinh^{2}{r}dt^{2}+dr^{2}\right)+\left(R_{0}\cosh{r}+\sqrt{R_{0}^{2}+l_{AdS}^{2}}\sinh{r}\cosh{t}\right)^{2}d\Omega_{d-1}^{2}\,$, (3.18)
where $R_{0}$ is the coordinate radius of the hole. For convenience, we can take the AdS radius to be $1$. The horizon is at $r=0$ similar to the spherical Rindler case. Taking the near horizon limit $r\rightarrow 0$, we end up with same metric (3.15):
$ds^{2}=-r^{2}dt^{2}+dr^{2}+R_{0}^{2}\,d\Omega_{d-1}^{2}$
in the leading order. So, in this geometry the flow will be same as given in (3.17). This flow reproduces the gravitational entropy of the hole, viz. ‘differential’ entropy. This method is only applicable in the near horizon limit of the hole. If one wants to construct a thread configuration, one needs to use the notion of covariant bit threads as formulated in [4].
4 First laws of entropy using bit-threads
The CPS formalism helps us to construct a representative flow for a valid bit thread configuration. So one can wonder if it is possible to represent different thermodynamical laws in terms of the thread picture. In this section, we will mainly focus on the first law of thermodynamics for the above mentioned entropies. Before proceeding further, recall that the symplectic current vanishes for Killing symmetries, i.e., $\boldsymbol{\omega}=0=\mathrm{d}\mathbf{k}_{\eta}$. So, we can construct a vector field from the codim-2 current as
$v^{a}_{th}=g^{ab}(\star\mathbf{k}_{\eta})_{b}\,$, (4.1)
which is similar to the map (2.13), and ensures that $v^{a}_{th}$ is divergenceless. We can also impose the norm bound as we did for $v^{a}$ in the previous section. Then we end up with a ‘thermodynamic variation flow’. This flow when integrated at the boundary of a Cauchy surface, results in variations of different thermodynamic quantities.
First law of BH thermodynamics
For a stationary, rotating, charge neutral, axially symmetric black hole with non-degenerate Killing vectors and a bifurcation surface, we derive the first law of the black hole thermodynamics using the covariant phase space formalism as briefly discussed in appendix B.3. For this kind of black holes, the Killing vectors associated with the horizon, time translation and rotation are related via the equation (B.21),
$\xi_{H}=\xi_{M}-\Omega_{H}\xi_{J},$
where, $\Omega_{H}$ represents the angular velocity of the horizon. Now, it is easy to find out the relation between the codim-2 currents corresponding to the Killing vectors and is given by,
$\mathbf{k}_{\rm{\eta_{S}}}=\frac{2\pi}{\kappa}\left(\mathbf{k}_{\rm{\eta_{M}}}-\Omega_{\rm{H}}\mathbf{k}_{\rm{\eta_{J}}}\right),$ (4.2)
where we have used the definitions (B.22) and relation (B.23) with $\kappa$ being the surface gravity of the horizon. For a stationary, axisymmetric black hole, these flows correspond to different conserved charges similar to the definitions (B.24):
$\delta M:=\oint_{\mathcal{S}_{\infty}}v_{th,M}\;;\quad\delta J:=\oint_{\mathcal{S}_{\infty}}v_{th,J}\;$; (4.3a)
$\delta S:=\oint_{\mathcal{B}}v_{th,S}\,.$ (4.3b)
Here we have used the notation (A.2). For simplicity we will omit the suffix ‘$th$’ hereafter. These fields also follow a similar relation like (4.2):
$v^{a}_{S}=\frac{2\pi}{\kappa}\left(v^{a}_{M}-\Omega_{H}v^{a}_{J}\right).$ (4.4)
Looking at the above equation carefully, it is evident that only one vector field is sufficient to construct the different conserved charges of a black hole. Now we will proceed the same way as we did while deriving the 1st law of BH thermodynamics. We start with the fact that for Killing vectors, the symplectic form vanishes,
$\int_{\Sigma}\boldsymbol{\omega}=0=\oint_{\partial\Sigma}\mathbf{k}_{\rm{\eta_{S}}}\,.$ (4.5)
Now to construct a perturbative vector field, we can use (4.1). Then the above equation becomes,
$\oint_{\partial\Sigma}\big{(}n_{a}v^{a}_{S}\big{)}\boldsymbol{\tilde{\epsilon}}=0\qquad\Rightarrow\qquad\int_{\Sigma}\big{(}\nabla_{a}v^{a}_{S}\big{)}\boldsymbol{\epsilon}=0.$ (4.6)
Here, $\boldsymbol{\tilde{\epsilon}}$ denotes the volume form of the boundary of the Cauchy surface $\partial\Sigma$ and $\boldsymbol{\epsilon}$ denotes the volume form of the Cauchy slice $\Sigma$. As the volume form $\boldsymbol{\epsilon}$ can vary according to the size of the black hole and other geometrical properties, the integrand should vanish. This gives us a ‘local’ 1st law of black hole thermodynamics as,
$\nabla_{a}v^{a}_{S}=0.$ (4.7)
So, the flow $v^{a}_{S}$ acts as a conserved current in the thermodynamics of black holes.
First law of cosmological horizons
The first law of thermodynamics for cosmological horizons (in static patch of dS) can also be expressed in terms of bit threads. For de Sitter space, the first law takes the form:
$T_{\rm{dS}}\delta S_{\rm{dS}}=\delta E,$ (4.8)
where $T_{\rm{dS}}=-\frac{1}{2\pi L}$ is the temperature and the minus sign reflects the fact that the cosmological horizon area decreases when energy is added to the system. In terms of perturbative bit threads, this can be written as:
$\nabla_{a}v_{S}^{a}=0,$ (4.9)
similar to the black hole case. This local law have few implications as follows:
The most direct interpretation is that certain concepts related to the thermodynamic quantities associated with the black hole, are locally conserved. This implies that changes in these quantities are not merely due to fluxes across the horizon but are governed by a local balance equation within the bulk itself. This provides a more granular understanding of how black hole (horizon) thermodynamics works.
First law of entanglement entropy
Here we will discuss how we can rewrite the first law of entanglement entropy in terms of bit threads. Like the previous section, we will use the canonical bit threads to construct a perturbative thread configuration in holographic spacetime. This approach was applied in [17], to reconstruct the metric and to show how linearised Einstein equation can be obtained from perturbative bit threads.
The first law of entanglement entropy says that the change in entanglement entropy is equal to the change in modular energy in first order. In [31], authors showed how one can extend the Iyer-Wald formalism in the subregion duality to reformulate the first law of entanglement entropy in a local way. The change in entanglement entropy and the modular energy can be written in terms of the codim-2 form $\mathbf{k}$ as follows [31]:
$\delta S_{A}=\int_{m(A)}\mathbf{k}_{\xi}\qquad\delta E_{A}=\int_{A}\mathbf{k}_{\xi},$ (4.10)
where, subscript $A$ denotes the boundary subregion for which we are looking at the EE $(S_{A})$, $m(A)$ is the minimal surface homologous to $A$ and $\delta E_{A}$ denotes the change in modular energy. The on-shell condition $\boldsymbol{\omega}=0=\mathrm{d}\mathbf{k}_{\xi}$ leads to the first law $\delta S_{A}=\delta E_{A}$ [^9] This can be easily checked as follows $\int_{\Sigma_{A}}\mathrm{d}\mathbf{k}_{\xi}=0=\int_{\partial\Sigma_{A}}\mathbf{k}_{\xi}\qquad\Rightarrow\int_{m(A)}\mathbf{k}_{\xi}=\int_{A}\mathbf{k}_{\xi}\qquad\Rightarrow\;\delta S_{A}=\delta E_{A}\,.$ . Now using the relation (4.1), we can rewrite the first law of entanglement entropy in terms of the bit thread vector fields,
$\int_{\Sigma_{A}}\big{(}\nabla_{a}v^{a}_{th}\big{)}\boldsymbol{\epsilon_{\Sigma_{A}}}=0\,$, (4.11)
where $\Sigma_{A}$ denotes the region in between $A$ and $m(A)$, and $\boldsymbol{\epsilon_{\Sigma_{A}}}$ is the volume form of that region. Now, by construction this perturbative bit thread is divergenceless, i.e., the integrand in the above equation does not depend on the size of the region $\Sigma_{A}$. So the first law of the entanglement entropy can be expressed in terms of the bit threads locally as:
$\nabla_{a}v^{a}_{th}=0.$ (4.12)
Here also thermodynamic flow acts like a conserved current ensuring that the EE and the modular energy are locally conserved.
First law of differential entropy
In a similar way, we can extend the above procedure for differential entropy. But there are few subtleties needed to be considered. Unlike the previous cases, the symplectic current does not vanish, which gives rise to a volume term in the first law of causal diamond [32]. Using the Complexity=Volume``Complexity\,=\,Volume" conjecture, the authors of [33] showed that we can obtain a first law of differential entropy in the holographic scenarios. So the perturbative flow corresponding to the differential entropy will no longer be divergenceless. Unlike the EE and Wald entropy cases, the diffeomorphism generator here is a conformal Killing vector which leads to the non-vanishing of the symplectic current. So, the fundamental identity (2.2) becomes,
$\boldsymbol{\omega}(\delta\Phi,\delta_{\chi}\Phi,\Phi)=\mathrm{d}\mathbf{k}_{\chi}(\delta\Phi,\Phi),$ (4.13)
where $\chi$ is the conformal Killing vector. The l.h.s. gives rise to the variation of volume of the spatial sphere (hole) whose causal diamond is under consideration and the r.h.s. gives rise to the variation of area of the boundary of that spatial sphere [32]. Integrating on that spatial spherical region $\sigma$,
$\int_{\sigma}\boldsymbol{\omega}(\delta\Phi,\delta_{\chi}\Phi,\Phi)=\int_{\sigma}\mathrm{d}\mathbf{k}_{\chi}(\delta\Phi,\Phi)=\int_{\partial\sigma}\mathbf{k}_{\chi}(\delta\Phi,\Phi).$ (4.14)
We can obtain a thermodynamic flow from $\mathbf{k}_{\chi}$, which is a perturbative Riemannian thread. On the other hand we can construct a perturbative Lorentzian thread configuration from the symplectic current using the map: $\delta\mathfrak{u}^{\mu}=g^{\mu\nu}(\ast\boldsymbol{\omega})_{n}$ as discussed in [34], where $\ast$ denotes the Hodge dual in a Lorentzian manifold. In terms of the threads, the above equation becomes,
$\int_{\sigma}\delta\mathfrak{u}=\int_{\partial\sigma}v_{th}=\int_{\sigma}\nabla_{a}v^{a}_{th}.$ (4.15)
This gives rise to the first law of differential entropy (or causal diamond). Locally we can write
$\delta\mathfrak{u}=\nabla_{a}v^{a}_{th}.$ (4.16)
The above equation gives us a relation between the Lorentzian and Riemannian (perturbative) bit threads.
5 Quantum corrections
We have mainly discussed the canonical bit threads in classical settings. In this section we will include the quantum corrections, then use the positivity of the relative entropy to obtain an energy condition which imposes some constraints on the quantum part. First we will briefly review the relative entropy and the quantum bit threads.
Relative entropy
In gravitational settings relative entropy is defined as [35],
$S_{\rm{rel}}\left(\rho_{A}^{\lambda}\parallel\rho_{A}^{0}\right)=S_{\rm{rel}}\left(g^{\lambda}\parallel g^{0}\right)=H_{\rm{bulk}}(g^{\lambda})-S_{\rm{HEE}}(g^{\lambda}).$ (5.1)
where, $H_{\rm{bulk}}$ is the bulk modular Hamiltonian or the gravitational energy, $\lambda$ denotes a one parameter family. Then the positivity of relative entropy tells that
$\frac{dS_{\rm{rel}}}{d\lambda}\Big{|}_{\lambda=0}\,\geq 0\,$,
which in gravitational terms can be written as,
$\dfrac{d}{d\lambda}\left[H_{\rm{bulk}}(g^{\lambda})-S_{\rm{HEE}}(g^{\lambda})\right]\geq 0\,.$ (5.2)
Again, by using the CPS formalism, one can rewrite the gravitational energy and entropy as [35],
$\dfrac{d}{d\lambda}H_{\rm{bulk}}(g^{\lambda})=\int_{A}\mathbf{k}_{\xi}~;\qquad\dfrac{d}{d\lambda}S_{\rm{HEE}}(g^{\lambda})=\int_{m(A)}\mathbf{k}_{\xi}\,.$ (5.3)
Therefore the positivity of relative entropy (5) can be expressed in terms of $\mathbf{k}$ as,
$\int_{A}\mathbf{k}_{\xi}-\int_{m(A)}\mathbf{k}_{\xi}\;\geq 0$
$\Rightarrow$~ $\int_{\Sigma_{A}}\mathrm{d}\mathbf{k}_{\xi}\;\geq 0.$ (5.4)
The boundary of the region $\Sigma_{A}$ is chosen to be: $\partial\Sigma_{A}=A-m(A)$. This relation will give us the energy condition by incorporating the idea of quantum bit threads.
Quantum bit threads
Quantum bit threads has been introduced recently in [11], and independently in [12]. Although both of them are closely related and can be shown to be the same in the leading orders, here the former one is discussed briefly. The quantum bit thread vector field $v$ is defined as $v=v_{0}+v_{q}$, such that:
$\int_{A}v=\int_{A}(v_{0}+v_{q})=S_{\rm{HEE}}^{0}+S_{\rm{bulk}},$ (5.5a)
but ,
$\int_{m(A)}v=\int_{m(A)}v_{0}=S_{\rm{HEE}}^{0}\,.$ (5.5b)
Here, $S_{\rm{HEE}}^{0}$ is the classical part of the HEE given by the area of the RT surface and $S_{\rm{bulk}}$ denotes the bulk entanglement entropy arises from quantum corrections. One can separate out the quantum part of $v$ as,
$\int_{\partial\Sigma_{A}}v_{q}=\int_{\Sigma_{A}}\nabla_{a}v_{q}^{a}=S_{\rm{bulk}}=\int_{\Sigma_{A}}s_{q}(x),$ (5.6)
with $\partial\Sigma_{A}=A-m(A)$. The integrand $s_{q}(x)$ was termed as entanglement contour in [11] but the authors emphasised that it need not to be a positive definite quantity as it was originally introduced in [36]. Later we will see that this quantity should be non-negative at least perturbatively in first order. So, to avoid any confusion, we can simply call the quantity $s_{q}(x)$ the ‘bulk entanglement density’ which gives rise to the bulk entanglement. The bulk entropy can also be written as the expectation value of the bulk modular Hamiltonian,
$S_{\rm{bulk}}=\Big{<}\hat{H}_{\rm{bulk}}\Big{>}=\int_{\Sigma_{A}}\Big{<}T^{\rm{bulk}}_{ab}(x)\Big{>}\xi^{a}\varepsilon^{b}.$
From the above definition it is clear that the divergenceless condition should be modified to incorporate the quantum part of the bit threads. Comparing the above two equations the modified divergence relation is given by,
$\nabla_{a}v_{q}^{a}=s_{q}(x)=\Big{<}T^{\rm{bulk}}_{ab}(x)\Big{>}\xi^{a}n^{b}.$ (5.7)
We will use this modified divergence relation along with the positivity of relative entropy to obtain an energy condition.
Quantum corrections and energy condition
The canonical bit threads prescription can be extended to include quantum corrections as well. First let us look at the classical case with matter. Recall the fundamental identity of the CPS formalism,
$\boldsymbol{\omega}(\delta\Phi,\mathcal{L}_{\xi}\Phi,\Phi)=\mathrm{d}\mathbf{k}_{\xi}+\xi^{a}\delta\boldsymbol{C}_{a},$ (5.8)
where, $\delta\boldsymbol{C}_{a}$ is related to the constraints of the theory. In presence of matter, $\delta\boldsymbol{C}_{a}=-\delta T_{ab}\varepsilon^{b}$. So even if the symplectic current vanishes (in the presence of a Killing vector $\xi$), $\mathbf{k}_{\xi}$ does not remain a closed form: $\mathrm{d}\mathbf{k}_{\xi}=-\xi^{a}\delta\boldsymbol{C}_{a}$. The vector $v^{a}_{th}$ constructed from this codim-2 form will also not be divergenceless101010 It implies that similar to the higher curvature gravity case [10], the divergence relation of bit threads should be modified in the presence of classical matter as well.. In the semiclassical regime, we consider the matter to be quantum fields. Then the constraint becomes $\delta\boldsymbol{C}_{a}=-\langle\delta T_{ab}\rangle\varepsilon^{b}$. For exact symmetries, the identity (5.8) reduces to,
$\mathrm{d}\mathbf{k}_{\xi}=-\xi^{a}\delta\boldsymbol{C}_{a}=\langle\delta T_{ab}\rangle\xi^{a}\varepsilon^{b}.$ (5.9)
Now, using the map (4.1), we can construct a vector field $\delta v^{a}$ which satisfies the following relation:
$\nabla_{a}\delta v^{a}=\langle\delta T_{ab}\rangle\xi^{a}\varepsilon^{b}.$ (5.10)
Here, $\delta v=\delta v_{0}+\delta v_{q}$. We have changed the notation from $v^{a}_{th}$ to $\delta v^{a}$ to emphasise that the quantum bit threads in discussion are perturbative in nature. The divergence is also related to the perturbed bulk entanglement density as: $\nabla_{a}\delta v^{a}=\delta s_{q}(x)$. Integrating both sides of the above equation over the region $\Sigma_{A}$, we find
$\int_{\Sigma_{A}}\nabla_{a}\delta v^{a}=\int_{\Sigma_{A}}\langle\delta T_{ab}\rangle\xi^{a}\varepsilon^{b}$
$\Rightarrow$ $\int_{A}\delta v=\int_{m(A)}\delta v+\int_{\Sigma_{A}}\langle\delta T_{ab}\rangle\xi^{a}\varepsilon^{b}$ (5.11)
This correctly captures the quantum corrections in this regime. Therefore the positivity inequality (5) of relative entropy can be rewritten in terms of bit threads vector field as follows
$\int_{\Sigma_{A}}\mathrm{d}\mathbf{k}_{\xi}\;\geq 0~~\Rightarrow~\int_{\Sigma_{A}}\nabla_{a}\delta v^{a}\;\geq 0\,.$ (5.12)
Using the modified divergence relation (5.10) we find,
$\ \int_{\Sigma_{A}}\delta s_{q}(x)\;\geq 0$
$\Rightarrow$ $\int_{\Sigma_{A}}\Big{<}\delta T^{\rm{bulk}}_{ab}(x)\Big{>}\xi^{a}\varepsilon^{b}\;\geq 0\,.$ (5.13)
This relation is true for any arbitrary size of the boundary region $A$. So the above inequality is true for the integrand itself which gives us the energy condition
$\Big{<}\delta T^{\rm{bulk}}_{ab}(x)\Big{>}\xi^{a}n^{b}\;\geq 0\,.$ (5.14)
We find that this energy condition is equivalent to the dominant energy condition (DEC)111111 To check that we use the physical form of the DEC for perturbative matter, $\delta T_{ab}\xi^{a}X^{b}\;\geq 0\,$, with $F^{a}=-\delta T^{a}_{\;\;b}\xi^{b}$ being a future pointing causal vector, and $\xi^{a}X_{a}<0$. For our case let $X^{a}=n^{a}$ is an unit normal timelike vector. Then it satisfies the condition $\xi^{a}n_{a}<0$. Also, as $F^{a}$ and $n^{a}$ both are future pointing and causal, they satisfy $F^{a}n_{a}\leq 0\qquad\Rightarrow\;-\delta T^{a}_{\;\;b}\xi^{b}n_{a}\leq 0\qquad\Rightarrow~~\delta T_{ab}\xi^{b}n^{a}\geq 0\,$, which is nothing but the classical version of the energy condition we have obtained. . So we conclude that if the perturbative bulk stress-energy tensor satisfies the DEC, then the bulk entanglement density should be positive. It subsequently implies the positivity of the relative entropy.
6 Bit threads in AdS-Vaidya
Till now we have been looking into static spacetimes except the spherical-Rindler case. But we took the near horizon limit (static in nature) which helped us to define a timelike Killing vector and use the covariant phase space formalism to obtain a divergenceless vector field. Then it was shown that this vector field can be thought as a bit thread configuration which gives us the differential entropy of a hole present in the spacetime. Now, we will apply the similar method to certain dynamical spacetimes.
Consider those dynamical spacetimes which can be mapped to static spacetimes via some conformal transformations. It has been showed that for such cases, the conformal relation plays an important role to derive the thermodynamics of the dynamical black hole [37]. Another example of this kind of geometry is the AdS-Vaidya spcaetime. Here, we will look into the AdS-Vaidya spacetime in 3 dimensions. The metric of this spacetime is
$ds^{2}=-\left(r^{2}-m(w)\right)dw^{2}+2dwdr+r^{2}dx^{2},$ (6.1)
where $m(w)$ is the mass function given by
$m(w)=M\,\theta(w)~~\rm{with}~~\text{$\theta(w)$}=\begin{cases}1,&\text{for $w\geq 0$}.\\ 0,&\text{for $w<0$}.\end{cases}$ (6.2)
Here we consider the mass $M$ to be used for parametric variations. The conformal Killing vector here is: $\xi=(\partial_{w},0,0)$. Using (2.20), we find
$16\pi G_{N}k^{wr}_{\xi}=-\frac{3\hat{\delta}M}{4\left(r^{2}-m(w)\right)^{2}}\left(r^{3}\theta(w)-M\delta_{DD}(w)\theta(w)\right)+\frac{\hat{\delta}M\delta_{DD}(w)}{r^{2}-m(w)}-\frac{\hat{\delta}M(2r\theta(w))}{r^{2}-m(w)},$
where $\delta_{DD}(.)$ represents the Dirac delta function. Here also we can find a perturbative flow by taking the Hodge dual of this form and contracting with the hypersurface metric. There exists a $\boldsymbol{\mathit{j}}_{\xi}$ such that $\mathbf{k}_{\xi}=\pm\hat{\delta}\boldsymbol{\mathit{j}}_{\xi}$ given by,
$16\pi G_{N}\mathit{j}^{wr}_{\xi}=$ $-\frac{3r^{2}\delta_{DD}(w)}{4m(w)-4r^{2}\theta(w)}-\frac{\delta_{DD}(w)\log\left(r^{2}-m(w)\right)}{4\theta(w)}$
$+\frac{r^{3}(11\theta(w)-8)-8r\left(r^{2}-m(w)\right)\log\left(r^{2}-m(w)\right)}{4\theta(w)\left(M-r^{2}\right)}\,.$ (6.3)
Now to find the vector $v^{a}_{cps}$, we need to specify a Cauchy surface. On the other hand, as the spacetime is dynamic, the extremal surface need not lie on a Cauchy surface. But to apply our method, we will restrict ourselves to a constant time slice and look into the spacelike extremal surface. We will take an approach similar to [38]. First, we rewrite the metric (6.1) in terms of $t$ and $r$ coordinates:
$ds^{2}=-f(r)dt^{2}+\dfrac{dr^{2}}{f(r)}+r^{2}dx^{2}\,$, (6.4)
where, $f(r)=r^{2}$ in AdS and $f(r)=(r^{2}-M)$ in BTZ. In the AdS phase, the current form (6) reduces to
$\mathit{j}^{wr}_{\xi}=\dfrac{1}{16\pi G_{N}}\dfrac{\mathfrak{C}}{r},$ (6.5)
where $\mathfrak{C}$ is an overall constant which can be fixed by imposing the boundary conditions. The above expression is quite similar to the pure AdS case which tells that the codim-2 current we obtained is correct. Using the map (2.13), we can find a divergenceless vector field,
$v^{a}_{cps}=\left(-v^{r}_{cps},v^{x}_{cps}\right)=-\dfrac{1}{4G_{N}}(\mathfrak{C}r,0).$ (6.6)
To check whether it gives rise to a flow or not depends on whether the vector field is proportional to the unit normal of the extremal surface corresponding to a boundary subregion. We can use the geodesic equation to find the minimal surface (spacelike geodesic) for a boundary subregion centred at $x=0$. The RT surface equation then reads:
$r=\dfrac{p_{x}}{\sqrt{x^{2}p_{x}^{2}-1}}\,$,
where $p_{x}$ is a conserved quantity along the RT surface. We can rearrange the above equation a little to obtain a nice form as,
$x^{2}+\dfrac{1}{r^{2}}=\dfrac{1}{p_{x}^{2}}\,.$ (6.7)
We can find the unit normal to this surface as given by,
$n^{a}=(n^{r},n^{x})=\left(-p_{x},\dfrac{xp_{x}}{r}\right).$ (6.8)
Now, comparing the vector (6.6) with the above unit normal of RT surface (6.7), one can easily check that $v^{a}_{cps}|_{RT}\neq\mathrm{C}n^{a}$. So, we do need to add $v^{a}_{\mathrm{d}}$ to make the vector field a bit thread flow. We solve the differential equation $\nabla_{a}v^{a}_{\mathrm{d}}=\nabla_{a}\nabla^{a}\varphi=0$ in the background (6.4) with proper boundary conditions, and find
$\varphi=\dfrac{1}{4G_{N}}\left[\dfrac{x}{p_{x}}\sqrt{r^{2}-p_{x}^{2}}\left(\mathfrak{C}r+p_{x}\right)+\dfrac{p_{x}}{r}-\dfrac{1}{2}\mathfrak{C}r^{2}x^{2}\right].$ (6.9)
The total vector field is given by,
$v^{a}=-\dfrac{1}{4G_{N}}\left(\mathfrak{C}r-p_{x}-\dfrac{\mathfrak{C}p_{x}}{\sqrt{x^{2}p_{x}^{2}-1}},\left(\mathfrak{C}r+p_{x}\right)\dfrac{\sqrt{r^{2}-p_{x}^{2}}}{r^{2}p_{x}}-\mathfrak{C}x\right).$ (6.10)
It reduces to the unit normal (6.8) at the RT surface (6.7), which confirms that the above vector field is a bit thread vector. One can also go to the Poincaré and find a flow, for which we have already shown that the vector field obtained from the CPS is not enough. We need to add the term $v^{a}_{\mathrm{d}}$ so that the total vector field becomes a bit thread configuration.
Now, we look into the BTZ phase ($w\geq 0$). In the $(t,r,x)$ coordinates, the form (6) becomes (3.3),
$j_{\xi}^{tr}=\dfrac{1}{16\pi G_{N}r}M,$
with AdS radius $L=1$. The vector obtain from this form is given by (3.4),
$v_{cps}^{a}=\left(-v_{cps}^{r},v_{cps}^{x}\right)=-\dfrac{1}{4G_{N}}\left(\dfrac{\sqrt{M}}{r}\sqrt{r^{2}-M},0\right).$ (6.11)
We need to look whether the vector field becomes proportional to the unit normal at the RT surface. We can find the RT surface equation using the geodesics entirely in the BTZ phase and using the metric (6.4). It is given by
$x=F\left(\sin^{-1}(r)\,\bigg{|}\,\frac{1}{p_{x}^{2}}\right)-E\left(\sin^{-1}(r)\,\bigg{|}\,\frac{1}{p_{x}^{2}}\right)-\dfrac{\sqrt{(r^{2}-p_{x}^{2})(r^{2}-M)}}{p_{x}\,r}\,$, (6.12)
where, $F(\mathsf{k}|\mathsf{m})$ is the incomplete elliptic integral of first kind and $E(\mathsf{k}|\mathsf{m})$ is the incomplete elliptic integral of second kind [^12] The incomplete elliptic integrals of first and second kind are given by[39], $F(\mathsf{k}|\mathsf{m})=\int_{0}^{\sin{\mathsf{k}}}\left[(1-t^{2})(1-\mathsf{m}\,t^{2})\right]^{-1/2}\,;\quad E(\mathsf{k}|\mathsf{m})=\int_{0}^{\sin{\mathsf{k}}}(1-t^{2})^{-1/2}(1-\mathsf{m}\,t^{2})^{1/2}.$ . The unit normal to this surface is given by,
$n^{a}=(n^{r},n^{x})=\dfrac{1}{r^{2}}\left(-rp_{x}\sqrt{r^{2}-M},\sqrt{r^{2}-p_{x}^{2}}\right).$ (6.13)
From the above expression and (6.11) it is clear that the resultant vector field $v^{a}_{cps}$ is not proportional to the normal $n^{a}$ at the RT surface (6.12). So, we need to add the ambiguity field $v_{\mathrm{d}}^{a}$ to make it a flow for a valid bit thread configuration. One can find this $v_{\mathrm{d}}^{a}$ by using PDE and the boundary conditions as we did for the pure AdS phase of the AdS-Vaidya geometry.
So we see that although the AdS3-Vaidya spacetime does not have a Killing vector, the presence of a conformal Killing vector can provide us the codim-2 forms and the corresponding vector field. This is surprising in a way that even in certain dynamical spacetimes, we can define (perturbative) bit thread flows using conformal Killing vectors rather than genuine Killing vectors. It extends the bit thread formalism beyond static spacetimes, though finding and interpreting the complete non-perturbative flow in general remains an open question.
7 Conclusions and Discussions
In this work, we have explored the the bit thread formalism and its broader applicability to geometric entropy concepts beyond the conventional HEE. Our study has centred on several key themes: formulating methods based on CPS formalism to obtain bit thread configurations, their connections to Wald and differential entropy, a reformulation of thermodynamic first laws in terms of flow dynamics, and preliminary steps toward accommodating quantum corrections and time dependence.
We attempt to construct bit thread configuration using quantities from the CPS formalism section. We found that although it naturally produces a divergenceless vector field $v_{cps}^{a}$, it generally does not meet the norm bound and saturation conditions required of HEE threads $v^{a}$. A gauge freedom $v_{\mathrm{d}}^{a}$ needed to satisfy the constraints of bit threads can be obtained by solving for harmonic function in $3$ dimensional spacetime. We showed that the bit threads vector field and $v^{a}_{cps}$ are in same equivalence class $v^{a}\sim v_{cps}^{a}$. This is one of the main results of our work. In spacetimes where the RT surface becomes horizon, the vector $v^{a}_{cps}$ is same as the bit thread. Then by applying suitable coordinate transformations to it results in a bit thread configuration in the original metric. We used this method to obtain a bit thread configuration for Poincaré AdS in section 2. This method does not require any prior knowledge of the RT surface and its unit normal, but surprisingly we find the resultant thread configuration to be same as the geodesic bit threads [18] up to an overall constant. These alternative ways for bit thread construction may be more suitable for certain setups.
We also extended this flow prescription to rewrite other entropies in the bit thread language. First we related the horizon entropy to this flow picture. In section 3, we showed how bit thread like vector can be found for AdS black holes which gives the Wald entropy once integrated over the bifurcation surface. Same can be done for non-AdS black holes also, as the Wald entropy is defined for any stationary black hole with bifurcation point. Similarly, the cosmological horizon yields a thread description. Another entropy which can be expressed in terms of threads is the differential entropy.
The CPS approach also inspired a new way to view thermodynamic first laws (Section 4). By utilising the Hodge dual of the current 2-form $\mathbf{k}_{\xi}$, we introduced the thermodynamic variation flow $v_{th}^{a}=g^{ab}(\star\mathbf{k}_{\xi})_{b}$, which captures the flux balances associated with entropy changes. These balance equations provide a geometric visualisation of how entropy variations arise from the bulk symplectic current, which is relevant for both black hole (and cosmological horizon) and entanglement first laws. For differential entropy, we find the first law relates the Lorentzian and Riemannian (perturbative) bit threads.
Finally, our preliminary exploration into quantum corrections (Section 5) and dynamical spacetimes (Section 6) may provide new directions for future work. Extending bit threads into the quantum regime requires modifying the norm bound or divergence conditions to account for bulk entanglement, potentially connecting to the generalised second law and quantum energy conditions. Time-dependent settings necessitate working with closed (d1)(d{-}1)-forms and pose significant challenges in constructing thread configurations in the absence of symmetries.
Here are some questions and future research directions:
  • Higher curvature gravity: The bit threads picture has been explored in higher curvature gravities in general [9], but an explicit expression for some particular theory is still missing. The method used here to construct a flow using the CPS formalism can be extended to these higher derivative gravities very naturally as the CPS formalism is applicable to more general formulations of gravity [15]. In [9], the authors showed how the norm bound changes for higher curvature gravity, while the vector field remains divergenceless. This property comes naturally following the CPS formalism and one should be able to construct a bit thread configuration by imposing the modified norm bound. It can be further extended to the cubic gravitational theories to possibly address the issue of ‘splitting’ as discussed in [40].
  • Towards covariant bit threads: The notion of covariant bit threads has been explored in [4]. Their exploration is mainly focused on the definition and general properties of the covariant bit threads. Unlike the Riemmanian ones, the covariant threads live in the whole spacetime and its flux is maximum at the HRT surface of the corresponding boundary region. The paper showed that we can find a covariant flow from a closed (d1)(d-1)-form obeying the required norm bound. If one wants to obtain these threads from CPS formalism, the Noether current (B.6) is a promising candidate which is a (d1)(d-1)-form residing in the spacetime unlike the codim-2 current discussed in this paper. Conservation ensures that the Noether current is a closed form (onshell). So one can find a divergenceless vector field by taking the Hodge dual of the Noether current,
    $V^{\mu}_{cov}=g^{\mu\nu}(\ast\boldsymbol{\mathrm{J}}_{\xi})_{\nu}\,.$ (7.1)
    We hope that this vector can be related to the ‘V-threads’ presented in [4] but the exact relation needs further investigation. Finding these threads explicitly will improve our understanding of the connection between entanglement and geometry. This method may also be helpful to extend the bulk reconstruction in the covariant setting similar to the Riemannian case as done in [17].
  • Quantum energy conditions: One interesting direction would be to derive the quantum focussing conjecture [41] and corresponding quantum null energy conditions [42] using this flow based prescription. It will also help to understand the covariant entropy bound and quantum Bousso bound through the bit threads perspective and may shed new light on the underlying structure of these quantities.
  • Generalised entanglement wedge: The construction of a flow from CPS formalism is also applicable to non-holographic spacetimes as discussed earlier. Recently, there has been a proposal of generalised entanglement wedge (GEW) in general gravitational spacetimes [43, 44], which states that for any gravitational bulk region $\mathtt{b}$, there is a GEW $\mathtt{E(b)}$ ($\supset\mathtt{b}$) on a static Cauchy surface. Then the fine grained entropy of the region $\mathtt{b}$ is given by the generalised entropy of the GEW $S_{gen}(\mathtt{E(b)})$. The authors of [45] extended the bit thread formulation for the GEW proposal in general gravitational spacetimes. The construction of bit threads using the CPS formalism provides a canonical way to extend the flow prescription to this GEW proposal.
  • Timelike entanglement entropy: It would be interesting to find a flow based picture for the timelike entanglement entropy [46]. One possible way is to use Wick rotated Euclidean metric in AdS and apply the method described in section 2. It may also be related to the pseudo entropy [47] which is more general than the timelike EE.
  • Connection to geometric modular flows: In CFTs, entanglement entropy of a subregion can be obtained by using the notion of modular geometric flows [48]. In the holographic context, the same can be obtained by using the bit threads. A possible connection has been already mentioned in [48]. The difference between these two formalism is that the geometric modular flow is uniquely determined whereas the configuration of bit threads of highly non-unique. Apparently, the bit threads and the geometric modular flows are normal to each other at the boundary subregion. If we consider the case in a Rindler wedge, the CPS method provides a canonical thread configuration. It will be worth exploring the connection between these two kind of flows in future.
Acknowledgements
We would like to thank Avik Chakraborty and Mrityunjay Nath for helpful discussions on various stages of this work. PD acknowledges the financial support by DST-INSPIRE fellowship (IF200253), Department of Science and Technology, Govt. of India. The work of MM is supported by DST-FIST initiative (SR/FST/LS-I/2020/621) and (GrantSR/FST/PSI-225/2016).
Appendix A Bit threads: a brief review
Here we will briefly review the notion of bit threads proposed by Headrick and Freedman [3]. We will start from the continuation version of well known max flow-min cut theorem originated in network theory.
Given an oriented $d$-dimensional Riemannian manifold with boundary, a ‘flow’ is a vector field $\upsilon$ satisfying,
$\nabla_{a}\upsilon^{a}=0\,,\qquad|\upsilon|\leq\mathrm{C}\,$, (A.1)
where, $\mathrm{C}$ is a positive quantity. A ‘surface’ is defined as a (d1)(d-1)-dimensional oriented submanifold. Then, flux of the flow through such a surface ‘$m$’ can be expressed as
$\int_{m}\upsilon:=\int_{m}\sqrt{\mathfrak{h}}\,n_{a}\upsilon^{a},$ (A.2)
where, $\mathfrak{h}$ is the determinant of the induced metric on $m$ and $n^{a}$ is the unit normal vector. The divergenceless condition (A.1) implies
$m\sim A\quad\Rightarrow\quad\int_{m}\upsilon=\int_{A}\upsilon\,$, (A.3)
where, $m\sim A$ denotes that the surfaces $m$ and $A$ are homologous to each other.
Figure 2: Schematic diagram of a Riemannian manifold showing a flow $\upsilon(A)$ and a surface $m(A)$ with minimal area, homologous to the boundary region $A$. The Riemannian MFMC tells that the flux of the flow (flow lines shown in brown) is proportional to the area of the surface $m(A)$ (shown in blue). This minimal surface acts as the bottleneck for the flow.
The norm bound $|\upsilon|\leq\mathrm{C}$ implies that the flux is bounded by the area of $m$:
$\int_{m}\upsilon\leq\mathrm{C}\int_{m}\sqrt{h}=\mathrm{C}\;{\cal{A}}(m)\,$, (A.4)
where ${\cal{A}}(m)$ denotes the area of $m$. Maximising on one side over all the flows $\upsilon$ and minimising on the other over all surfaces homologous to $A$ yields
$\max_{\upsilon}\int_{A}\upsilon\leq\mathrm{C}\min_{m\sim A}\;{\cal{A}}(m)\,.$ (A.5)
The Riemannian max-flow min-cut (MFMC) theorem [3, 49] states that the above inequality is saturated:
$\max_{\upsilon}\int_{A}\upsilon=\mathrm{C}\min_{m\sim A}\;{\cal{A}}(m)\,.$ (A.6)
The inequalities (A.4) provide an an upper bound for the flux of any flow via the area of a minimal surface $m(A)$ (homologous to $A$), which is the bottleneck (as shown in fig. 2) [^13] We thank Ayushi Suman for the figure.. The MFMC theorem (A.6) implies that $n_{a}\upsilon^{a}=\mathrm{C}$, and hence $\upsilon^{a}=\mathrm{C}n^{a}$, everywhere on $m(A)$.
In the holographic context, one can use the above MFMC theorem to rewrite the RT formula for entanglement entropy. In a bulk theory, the Riemannian manifold is a constant-time slice, and $A$ is a boundary subregion in conformal field theory. Then, with the help of equation (A.6), and by setting $\mathrm{C}=1/4G_{N}$, the RT formula can be reexpressed as,
$S_{\rm{HEE}}(A)=\max_{\upsilon}\int_{A}\upsilon\,.$ (A.7)
Appendix B Covariant phase space formalism
The covariant phase space formalism [50, 15, 16] is an approach to study Hamiltonian mechanics in which one constructs a phase space by taking the set of solutions of the equation of motions instead of the generalised coordinates and their conjugates. To study the entropy and thermodynamics of black holes, people mainly use the Iyer-Wald formalism [15]. There have been many modifications and generalisations to this formalism. One such example is [24], where the authors considered the boundary Lagrangian separately. However, the entropy obtained from both of these formalisms is same. Here, we will briefly discuss both of these formalisms, the Wald entropy and the well known first law of black hole thermodynamics.
B.1 Iyer-Wald Formalism
In CPS, the phase space is considered to be a manifold $\mathcal{M}$ with a symplectic two-form $\Omega$. Then from the symplectic form, one can define the Hamiltonian. The action for a generally invariant theory is
$\mathrm{S}[\Phi]=\int_{\mathcal{M}}\boldsymbol{\mathrm{L}}(\Phi),$ (B.1)
where, $\boldsymbol{\mathrm{L}}$ is the Lagrangian d-form for any d-dimensional theory, and $\Phi$ denotes all the dynamical fields collectively. Taking variation of this action leads to
$\delta\boldsymbol{\mathrm{L}}[\Phi]=\boldsymbol{\mathrm{E}}_{\Phi}\delta\Phi+\mathrm{d}\mathbf{\Theta}_{\rm{IW}}\left(\delta\Phi,\Phi\right),$ (B.2)
where $\delta\Phi$ represents a generic field perturbation and provides the basis of (tangent space of) the phase space $\mathcal{P}$. The equations of motion are given by the set of equations $\boldsymbol{\mathrm{E}}_{\Phi}=0$. The quantity $\mathbf{\Theta}_{\rm{IW}}$ is called the symplectic potential. It is a (d1)(d-1)-form in spacetime $\mathcal{M}$ and one form on the phase space, i.e., $\mathbf{\Theta}_{\rm{IW}}$ is a (d1;1)(d-1;1)-form. Then one can define the quantity called (pre)symplectic current as
$\boldsymbol{\omega}_{\rm{IW}}(\delta_{1}\Phi,\delta_{2}\Phi,\Phi)=\delta_{1}\mathbf{\Theta}_{\rm{IW}}(\delta_{2}\Phi,\Phi)-\delta_{2}\mathbf{\Theta}_{\rm{IW}}(\delta_{1}\Phi,\Phi).$ (B.3)
Here, $\delta_{1}\Phi$ and $\delta_{2}\Phi$ are two arbitrary field perturbations, $\mathrm{d}$’ denotes the exterior derivative over the spacetime $\mathcal{M}$ and $\delta$’ represents the exterior derivative over the phase space manifold $\mathcal{P}$. Using this notation, it is evident that $\boldsymbol{\omega}_{\rm{IW}}$ is a two-form on the phase space, or more compactly, is a (d1;2)(d-1;2)-form. A symplectic form then can be constructed by integrating the (pre) symplectic current over a codimension-one spacelike surface $\Sigma$,
$\mathbf{\Omega}_{\rm{IW}}(\delta_{1}\Phi,\delta_{2}\Phi,\Phi)=\int_{\Sigma}\boldsymbol{\omega}_{\rm{IW}}(\delta_{1}\Phi,\delta_{2}\Phi,\Phi).$ (B.4)
This $\mathbf{\Omega}_{\rm{IW}}$ is non-degenerate, and it does not depend on the choice of the Cauchy surface $\Sigma$. If $\mathcal{F}$ is the set of solutions of the equations of motion, then $(\mathcal{F};\mathbf{\Omega})$ constitutes a well-defined phase space denoted by $\cal{P}$}.
Now, given a transformation generated by $\epsilon$, one can define different quantities like Noether charge, Noether current, and the Hamiltonian (or, more generally a conserved charge). For diffeomorphism invariant theories, covariance tells that $\delta_{\xi}\Phi=\mathcal{L}_{\xi}\Phi$, where $\xi$ is the diffeomorphism generator. So the (pre)-symplectic current becomes,
$\boldsymbol{\omega}_{\rm{IW}}(\delta\Phi,\mathcal{L}_{\xi}\Phi,\Phi)=\delta\mathbf{\Theta}_{\rm{IW}}(\mathcal{L}_{\xi}\Phi,\Phi)-\mathcal{L}_{\xi}\mathbf{\Theta}_{\rm{IW}}(\delta\Phi,\Phi).$ (B.5)
One can now introduce the Noether current $\boldsymbol{\mathrm{J}}$ as,
$\boldsymbol{\mathrm{J}}_{\xi}\equiv\mathbf{\Theta}_{\rm{IW}}(\mathcal{L}_{\xi}\Phi,\Phi)-\xi.\boldsymbol{\mathrm{L}}~.$ (B.6)
A simple check shows that this current is a closed form, and locally, it can be made an exact form, which leads to the definition of the Noether charge,
$\boldsymbol{\mathrm{J}}_{\xi}=\mathrm{d}\boldsymbol{\mathrm{Q}}_{\xi}.$ (B.7)
The Hamiltonian can then be written as,
$\delta H_{\xi}=\mathbf{\Omega}_{\rm{IW}}(\delta\Phi,\mathcal{L}_{\xi}\Phi,\Phi)=\int_{\Sigma}\boldsymbol{\omega}_{\rm{IW}}.$ (B.8)
The above expression can be written in terms of the Noether current and Noether charge as well. Now, we will talk about the recent modified version of the covariant phase space with boundaries.
B.2 CPS with Boundaries
Recently Harlow and Wu generalised the CPS formalism by adding a boundary term and the result is the covariant phase space formalism with boundaries (CPSB) [24]. In this kind of formulation, the action is written as,
$\mathrm{S}[\Phi]=\int_{\mathcal{M}}\boldsymbol{\mathrm{L}}\big{(}\Phi,\Xi\big{)}+\int_{\partial\mathcal{M}}\boldsymbol{{\ell}}\big{(}\Phi,\Xi\big{)}$ (B.9)
with boundary conditions for $\Phi^{i}(x)$ on the spatial boundary $\Gamma$. Here, $\Phi$ is dynamical and $\Xi$ is background field. Then one imposes that the action is well posed, i.e., $\Phi$ obeys EOM iff $\mathrm{S}$ is stationary upto future/past terms at $\Phi$. Similar to IW formalism, here also the Lagrangian is considered to be a $d$-form (for $d$-dimensional spacetime/manifold) instead of a scalar. Therefore, the variation of the Lagrangian can be written as,
$\delta\boldsymbol{\mathrm{L}}=\boldsymbol{\mathrm{E}}_{i}\delta\Phi^{i}+\mathrm{d}\mathbf{\Theta}_{\rm{HW}}$ (B.10)
where, $\mathrm{d}$’ is exterior derivative on space-time, $\mathbf{\Theta}_{\rm{HW}}$ is called symplectic potential. Thus, the variation of the action is given by
$\delta\mathrm{S}=\int_{\mathcal{M}}\boldsymbol{\mathrm{E}}_{i}\delta\Phi^{i}+\int_{\Sigma_{\pm}}\big{(}\mathbf{\Theta}_{\rm{HW}}+\delta\boldsymbol{\ell}\big{)}+\int_{\Gamma}\big{(}\mathbf{\Theta}_{\rm{HW}}+\delta\boldsymbol{\ell}\big{)}$ (B.11)
As mentioned earlier, the action is required to be stationary. The stationarity of $\mathrm{S}$ implies $\delta\mathrm{S}=0$. This condition results to two important equations:
$\boldsymbol{\mathrm{E}}_{i}=0\quad(Equations\;of\;Motion)$ (B.12)
and, the other one is the non-trivial one,
$\big{(}\mathbf{\Theta}_{\rm{HW}}+\delta\boldsymbol{\ell}\big{)}\mid_{\Gamma}=\mathrm{d}\boldsymbol{\mathrm{C}}$ (B.13)
for any (d2)(d-2)-form $\boldsymbol{\mathrm{C}}$. Now, one can define a quantity named pre-symplectic current, $\boldsymbol{\omega}_{\rm{HW}}\equiv\delta(\mathbf{\Theta}_{\rm{HW}}-\mathrm{d}\boldsymbol{\mathrm{C}})\mid_{\tilde{\mathcal{P}}}$ with certain properties. Integrating the pre-symplectic current over a Cauchy slice one obtains the pre-symplectic form,
$\tilde{\mathbf{\Omega}}_{\rm{HW}}=\int_{\Sigma}\boldsymbol{\omega}_{\rm{HW}}$ (B.14)
[…]
$\delta H_{\xi}=-X_{\xi}.\tilde{\mathbf{\Omega}}_{\rm{HW}},$ (B.15)
where, $X_{\xi}\equiv\int d^{d}\mathrm{x}\;\mathcal{L}_{\xi}\;\Phi^{i}(x)\dfrac{\delta}{\delta\Phi^{i}(x)}$. Noether current is defined as,
$\boldsymbol{\mathrm{J}}_{\xi}\equiv X_{\xi}.\mathbf{\Theta}_{\rm{HW}}-\xi.\boldsymbol{\mathrm{L}}.$ (B.16)
Simple calculation using the above three equations gives,
$X_{\xi}.\tilde{\mathbf{\Omega}}_{\rm{HW}}=\delta\left(\int_{\Sigma}\boldsymbol{\mathrm{J}}_{\xi}+\int_{\partial\Sigma}\big{(}\xi.\boldsymbol{{\ell}}-X_{\xi}.\boldsymbol{\mathrm{C}}\big{)}\right).$ (B.17)
Evidently, from equations (B.15) and (B.17), one obtains the Hamiltonian,
$H_{\xi}=\int_{\Sigma}\boldsymbol{\mathrm{J}}_{\xi}+\int_{\partial\Sigma}\big{(}\xi.\boldsymbol{{\ell}}-X_{\xi}.\boldsymbol{\mathrm{C}}\big{)}.$ (B.18)
This is true for any kind of higher derivative theory. Some examples can be seen in [24]. There are some special theories (like the theory of general relativity), where $\boldsymbol{\mathrm{L}}$ is covariant under arbitrary $\xi^{\mu}$. This gives us $\mathrm{d}\boldsymbol{\mathrm{J}}_{\xi}=0$ and $\boldsymbol{\mathrm{J}}_{\xi}$ becomes an exact form. So, one can define a (d2)(d-2)-form $\boldsymbol{\mathrm{Q}}_{\xi}$ such that,
$\boldsymbol{\mathrm{J}}_{\xi}=\mathrm{d}\boldsymbol{\mathrm{Q}}_{\xi},$ (B.19)
$\boldsymbol{\mathrm{Q}}_{\xi}$ being the “Noether charge”. For these special theories, the Hamiltonian gets simplified to,
$H_{\xi}=\int_{\partial\Sigma}\left(\boldsymbol{\mathrm{Q}}_{\xi}+\xi.\boldsymbol{{\ell}}-X_{\xi}.\boldsymbol{\mathrm{C}}\right),$ (B.20)
which shows that the Hamiltonian is a totally boundary term.
B.3 Black Hole Thermodynamics
Consider a stationary rotating black hole with non-degenerate Killing vectors. Further, assume the black hole is axially symmetric, has a bifurcation, and has no electric charge. The horizon Killing vector is denoted by $\xi_{H}$. The Killing vector corresponding to the time evolution is $\xi_{M}$, and the Killing vectors corresponding to the rotation is $\xi_{J}$. These Killing vectors are related by,
$\xi_{H}=\xi_{M}-\Omega_{H}\xi_{J},$ (B.21)
where, $\Omega_{H}$ represents the angular velocity of the horizon. Redefine the generators
$\eta_{M}=\xi_{M};\qquad\eta_{J}=\xi_{J};\qquad\eta_{S}=\frac{2\pi}{\kappa}\xi_{H},$ (B.22)
where, $\kappa$ is the surface gravity of the black hole. The obvious relation between these new generators is given by,
$\eta_{S}=\frac{2\pi}{\kappa}\left(\eta_{M}-\Omega_{H}\eta_{J}\right)$ (B.23)
Now, one can define the conserved charges corresponding to the generators. The conserved charges for time evolution and rotation are defined by integrating the codimensional-2 currents $\mathbf{k}_{\eta_{M}}$ and $\mathbf{k}_{\eta_{J_{i}}}$ at the spatial infinity,
$\delta M:=\oint_{\mathcal{S}_{\infty}}\mathbf{k}_{\eta_{M}}(\delta\Phi,\Phi);\\$ (B.24a)
$\delta J:=\oint_{\mathcal{S}_{\infty}}\mathbf{k}_{\eta_{J}}(\delta\Phi,\Phi),$ (B.24b)
and the conserved charge corresponding to the horizon Killing vector by integrating $\mathbf{k}_{\eta_{S}}$ at the bifurcation surface,
$\delta S:=\oint_{\mathcal{B}}\mathbf{k}_{\eta_{S}}(\delta\Phi,\Phi).$ (B.24c)
These conserved charges $\delta M$, $\delta J$, and $\delta S$ are the mass of the black hole, the angular momentum of the black hole horizon, and the entropy of the black hole, respectively. To obtain the first law of black hole thermodynamics using the conserved charges and their corresponding codim-2 currents, recall the fundamental identity (2.2),
$\int_{\Sigma}\boldsymbol{\omega}(\delta\Phi,\mathcal{L}_{\xi_{\rm{H}}}\Phi,\Phi)=\int_{\Sigma}\mathrm{d}\mathbf{k}_{\eta}(\delta\Phi,\Phi).$ (B.25)
The Cauchy surface $\Sigma$ has boundary $\partial\Sigma=\mathcal{S}_{\infty}\cup(-\mathcal{B})$, where $\mathcal{B}$ is a surface (here, the bifurcation surface) homologous to $\mathcal{S}_{\infty}$ (the spatial infinity). As $\xi_{H}$ is a Killing vector, any variation due to this should vanish, i.e., $\mathcal{L}_{\xi_{H}}\Phi=0$, or more preciously,
$\int_{\Sigma}\boldsymbol{\omega}(\delta\Phi,\mathcal{L}_{\xi_{\rm{H}}}\Phi,\Phi)=0,$ (B.26)
which immediately leads to,
$\oint_{\mathcal{S}_{\infty}}\mathbf{k}_{\rm{\eta_{M}}}(\delta\Phi,\Phi)-\Omega_{\rm{H}}\oint_{\mathcal{S}_{\infty}}\mathbf{k}_{\rm{\eta_{J}}}(\delta\Phi,\Phi)-\frac{\kappa}{2\pi}\oint_{\mathcal{B}}\mathbf{k}_{\rm{\eta_{S}}}(\delta\Phi,\Phi)=0,$ (B.27)
where the relation (4.2) has been used. Now, using the definitions of the conserved charges (B.24),
$\delta M-\Omega_{\rm{H}}\delta J-\frac{\kappa}{2\pi}\delta S=0.$
The temperature of the horizon is defined as $T_{\rm{H}}=\frac{\kappa}{2\pi}$. Substituting it, one obtains the well known 1st law of black hole thermodynamics,
$\delta M=T_{\rm{H}}\delta S+\Omega_{\rm{H}}\delta J.$ (B.28)
Appendix C Flow in D11-brane background
$D1$-brane is an object in $10D$ supergravity. An effective theory for this system can be obtained in 3 dimensions by compactifying it on a 7-sphere. This effective $3D$ Einstein-dilaton system is governed by the action [51],
$\mathrm{S}_{\rm{ED}}=\dfrac{1}{16\pi G_{N}}\int d^{3}x\sqrt{-g}\left[R-\dfrac{\nu}{2}\partial_{\mu}\varphi\partial^{\mu}\varphi-{\cal{V}}(\varphi)\right],$ (C.1)
where, $\nu=\frac{16}{9}$, $\varphi$ is the dilaton and its potential ${\cal{V}}(\varphi)=-\frac{24}{L^{2}}e^{\frac{4}{3}\varphi}$. The $3D$ metric of the $D1$-brane is given by,
$ds^{2}_{\rm{D1}}=-\mathcal{C}_{T}(r)^{2}dt^{2}+\mathcal{C}_{R}(r)^{2}dr^{2}+\mathcal{C}_{X}(r)^{2}dx^{2},$ (C.2)
with the metric components,
$\mathcal{C}_{T}(r)^{2}=\left(\frac{r}{L}\right)^{8}f~;~\mathcal{C}_{R}(r)^{2}=\frac{1}{f}\left(\frac{r}{L}\right)^{2}~;~\mathcal{C}_{X}(r)^{2}=\left(\frac{r}{L}\right)^{8},$ (C.3)
where $f=1-\dfrac{r_{0}^{6}}{r^{6}}$, $r_{0}$ being the non-extremal parameter. And the dilaton is given by,
$\varphi=-3\ln{\left(\dfrac{r}{L}\right)}.$ (C.4)
The metric (C.2) contains two constants: the AdS radius $L$ and the non-extremal parameter $r_{0}$. We can use either of these two or both to do the parametric variations as discussed in section 2. For simplicity we will use the AdS radius $L$ for the variatio. Unlike previous examples, this solution consists of both the gravity and a scalar (dilaton) sector. For the gravity part we can use (2.20) to find the non-zero component of the codim-2 current:
$k^{tr}_{EH}=-\frac{L\left(r^{6}+2r_{0}^{6}\right)}{4\pi G_{N}r^{9}}\hat{\delta}L,$ (C.5)
where the suffix $EH$ denotes the Einstein-Hilbert part of the action. For the scalar sector we can use the expression [23],
$k^{\alpha\beta}_{Sc}=\dfrac{1}{8\pi G_{N}}\left[\xi^{\beta}\tilde{\nu}\nabla^{\alpha}\phi\delta\phi-(\alpha\leftrightarrow\beta)\right],$ (C.6)
with $\tilde{\nu}=\frac{\nu}{4}=\frac{4}{9}$. Non-zero component of this current is given by,
$k^{tr}_{Sc}=\frac{L}{4\pi G_{N}r^{3}}\left(1-\frac{r^{6}_{0}}{r^{6}}\right)\hat{\delta}L.$ (C.7)
By adding these two different contributions we get the total codim-2 current,
$k^{ab}$ $=k^{ab}_{EH}+k^{ab}_{Sc}$
$\Rightarrow k^{tr}$ $=-\frac{3L^{2}r^{6}_{0}}{4\pi G_{N}r^{9}}\hat{\delta}L.$ (C.8)
Now, we can find a (d2)(d-2)-form $\boldsymbol{\mathit{j}}_{\xi}$ such that $\mathbf{k}=-\hat{\delta}\boldsymbol{\mathit{j}}$ given as,
$j^{tr}=\dfrac{1}{4\pi G_{N}}\left(\dfrac{r_{0}^{6}L^{3}}{r^{9}}\right).$ (C.9)
Restricting this form on a constant time slice and taking the Hodge dual, we find a divergenceless vector field
$v^{a}_{cps}=\left(-v^{r}_{cps},v^{x}_{cps}\right)=-\dfrac{1}{4G_{N}}\left(\mathfrak{E}\dfrac{r_{0}^{4}L}{r^{5}}\sqrt{1-\dfrac{r_{0}^{6}}{r^{6}}},0\right).$ (C.10)
Now, to check if it gives rise to bit threads, we need to see whether this vector is proportional to the unit normal of a RT surface corresponding to a boundary subregion. Let us first try to find the RT surface equation. Finding this surface for an interval (in the boundary) is akin to going away from the centre of the bulk, i.e., taking the limit $r\rightarrow\infty$ (or $r\gg r_{0}$). In this limit the metric (C.2) reduces to
$ds^{2}_{\rm{D1}}|_{r\gg r_{0}}=-\left(\frac{r}{L}\right)^{8}dt^{2}+\left(\frac{r}{L}\right)^{2}dr^{2}+\left(\frac{r}{L}\right)^{8}dx^{2}\,.$ (C.11)
We can try to further simplify it. Substituting $y=r^{4}$ yields
$ds^{2}=y^{2}(-dt^{2}+dx^{2})+\dfrac{dy^{2}}{16y}.$ (C.12)
RT surface can be obtained by integrating over the following differential equation,
$dx=\dfrac{y_{0}dy}{4y^{3/2}\sqrt{y^{2}-y^{2}_{0}}}\,$, (C.13)
where, $y_{0}$ is the deepest point of the RT surface in the bulk. Integrating both sides of the above equation gives us the RT surface equation
$4y^{2}_{0}x^{2}y=(y^{2}-y_{0}^{2})\,{{}_{2}}F_{1}\left(\,\dfrac{1}{4}\,,\,1\,;\,\dfrac{3}{4}\,;\,\dfrac{y^{2}}{y_{0}^{2}}\,\right)^{2}.$ (C.14)
We can find the unit normal to this surface by using (C.13), given by
$n^{a}=(n^{y},n^{x})=\dfrac{1}{y^{2}\sqrt{y^{2}-y^{2}_{0}}}\left(-4y_{0}y^{3/2},\sqrt{y^{2}-y^{2}_{0}}\right).$ (C.15)
[…]
$v^{a}_{cps}=\left(-v^{y}_{cps},v^{x}_{cps}\right)=-\dfrac{1}{4G_{N}}\left(\mathfrak{K}\sqrt{y},0\right).$ (C.16)
This vector field also does not matches with the unit normal (C.15) at the RT surface. So, we need to add a non-trivial vector field $v^{a}_{\mathrm{d}}$ or the harmonic function $\varphi$. Solving the Laplace’s equation in the background (C.12) and using proper boundary conditions, we find
$\varphi=\dfrac{1}{4G_{N}}\left(x-\dfrac{1}{2\sqrt{y_{0}}}E\left(\dfrac{1}{2}\sec^{-1}(\dfrac{y}{y_{0}})\,\bigg{|}\,\frac{1}{p_{x}^{2}}\right)-\dfrac{\mathfrak{K}}{32}\sqrt{y}-\dfrac{4y_{0}y^{3/2}}{(y^{2}-y_{0}^{2})^{3/2}}\left(\dfrac{x^{2}}{2}-x\mathfrak{f}(y)\right)\right),$ (C.17)
with $\mathfrak{f}(y)=\frac{\sqrt{y^{2}-y_{0}^{2}}}{2y_{0}\sqrt{y}}{{}_{2}}F_{1}\left(\,\frac{1}{4}\,,\,1\,;\,\frac{3}{4}\,;\,\frac{y^{2}}{y_{0}^{2}}\,\right)$.
Appendix D Bit threads from PDE
The max-flow formulation of holographic entanglement entropy, seeking a divergenceless vector field $v^{a}$ with bounded norm $|v|\leq\mathrm{C}$ that maximises flux $\int_{A}v$, bears a structural resemblance to certain problems involving partial differential equations (PDEs). This appendix aims to briefly outline these connections.
One relevant connection arises from potential flow theory. If the bit thread field $v^{a}$ were derivable from a scalar potential $\psi$ (i.e., $v^{a}=\nabla^{a}\psi$), the divergenceless condition $\nabla_{a}v^{a}=0$ would translate into the Laplace equation $\nabla^{2}\psi=0$. The problem would then involve solving Laplace equation subject to the non-linear norm constraint $|\nabla\psi|\leq\mathrm{C}$ and appropriate boundary conditions related to the flux maximisation for region $A$.
Figure 3: Flow lines for bit threads obtained from solving the PDE: $\nabla_{a}v_{PDE}^{a}=0$ and imposing the conditions $v_{PDE}^{a}|_{{RT}}=n^{a}$ and $|v_{PDE}|\leq 1$. The blue horizontal line denotes the boundary subregion and the red semicircle is the corresponding RT surface.
Here we construct a flow directly from the PDE: $\nabla_{a}v_{PDE}^{a}=0$ and using the method of characteristics, we find a general solution. Then imposing the conditions $v_{PDE}^{a}|_{\footnotesize{RT}}=n^{a}/4G_{N}$ and $|v_{PDE}|\leq 1/4G_{N}$. We have not shown for the scalar $\psi$, as we are interested in the vector $\nabla^{a}\psi$ itself. For convenience, we solved the PDE in the Poincaré AdS3 background (2.19) and find a vector field given by,
$v_{PDE}^{a}=\dfrac{1}{4G_{N}}\left(\frac{z^{2}\left(R\sqrt{R^{2}-x^{2}}-Rz\right)}{\left(R^{2}-x^{2}\right)^{3/2}+x^{2}}+\frac{z^{2}}{R},\frac{xz}{R}\right).$ (D.1)
This is a valid bit thread configuration. Note that this vector is only well defined from the boundary subregion $A$ to the corresponding RT surface $m(A)$. Beyond that we need to glue some other thread configuration (like the one in section 2 or the geodesic ones) such that $v^{a}$ remains continuous and differentiable at the RT surface. A similar method was used for ‘maximally packed flows’ in [18]. The flow lines of $v_{PDE}^{a}$ are shown in figure 3. For simple geometries this method can also be applied to construct a flow for bit threads.
Appendix E Coordinate systems of AdS3
Here we will discuss the relation between two different coordinate patches of the AdS spacetime, namely the Poincaré and Rindler AdS. We will mainly follow the notations of [27]. The geometry of AdSd+1 can be expressed by embedding it to a $\mathbb{R}^{2,d}$ space as
$Y_{-1}^{2}-Y^{2}_{0}+Y_{1}^{2}+\cdots+Y_{d}^{2}=-L^{2},$ (E.1)
with metric
$ds^{2}=-dY_{-1}^{2}-dY^{2}_{0}+dY_{1}^{2}+\cdots+dY_{d}^{2}.$ (E.2)
As discussed in section 2, we are mainly interested in 3 dimensional AdS spacetime. So, the above geometry reduces to AdS3 embedded in $\mathbb{R}^{2,2}$ space as
$Y_{-1}^{2}-Y^{2}_{0}+Y_{1}^{2}+Y_{2}^{2}=-L^{2},$ (E.3)
with metric
$ds^{2}=-dY_{-1}^{2}-dY^{2}_{0}+dY_{1}^{2}+dY_{2}^{2}.$ (E.4)
This geometry is connected to the Poincaré AdS3 via
$Y_{-1}+Y_{2}=\dfrac{L^{2}}{z},\qquad Y^{\mu}=\dfrac{L}{z}x^{\mu}\quad\mathrm{with}~\;\mu=0,1.$ (E.5)
Combing the above equation with the hyperboloid (E.3) gives the standard Poincaré AdS3 metric
$ds^{2}_{P}=\dfrac{L^{2}}{z^{2}}\left(-dt^{2}+dx^{2}+dz^{2}\right),$ (E.6)
where, $t$ is the time coordinate in Poincaré patch.
The Rindler-AdS space is another foliation of AdS3 we are interested in, given by
$Y_{-1}=\rho\cosh{u},\quad Y_{0}=\tilde{\rho}\sinh{\left(\frac{\tau}{L}\right)},\quad Y_{2}=\tilde{\rho}\cosh{\left(\frac{\tau}{L}\right)},\quad Y_{1}=\rho\sinh{u}.$ (E.7)
Substituting these in (E.3) yields $\tilde{\rho}=\sqrt{\rho^{2}-L^{2}}$ and the metric becomes
$ds^{2}_{R}=-\left(\dfrac{\tilde{\rho}^{2}}{L^{2}}\right)d\tau^{2}+\left(\dfrac{\tilde{\rho}^{2}}{L^{2}}\right)^{-1}d\rho^{2}+\rho^{2}du^{2}.$ (E.8)
[…]
$Y_{-1}+Y_{2}=\dfrac{L^{2}}{z}=\rho\cosh{u}+\tilde{\rho}\cosh{\left(\frac{\tau}{L}\right)},\quad Y_{0}=\dfrac{L}{z}t=\tilde{\rho}\sinh{\left(\frac{\tau}{L}\right)},\quad Y_{1}=\dfrac{L}{z}x=\rho\sinh{u}.$ (E.9)
As our main objective is to use the coordinate transformations to find a vector field in terms of one coordinate system from the other on a fixed Cauchy surface, we will look at the above relations with $t=0=\tau$ which is a common Cauchy surface in both the coordinate systems. Now, the above relations simplify to
$Y_{-1}+Y_{2}=\dfrac{L^{2}}{z}=\rho\cosh{u}+\tilde{\rho},\qquad Y_{1}=\dfrac{L}{z}x=\rho\sinh{u}.$ (E.10)
From these two equations, we find the coordinate transformations given by
$\rho=\dfrac{1}{2z}\sqrt{\left(L^{2}+x^{2}+z^{2}\right)^{2}-4L^{2}x^{2}},$ (E.11a)
$u=\operatorname{arctanh}{\left(\dfrac{2Lx}{L^{2}+x^{2}+z^{2}}\right)}.$ (E.11b)
Similarly, the inverse coordinate transformations are
$z=\dfrac{L^{2}}{\rho\cosh{u}+\tilde{\rho}}\,,\qquad x=\dfrac{L\rho\sinh{u}}{\rho\cosh{u}+\tilde{\rho}}$. (E.12)
Using these coordinate transformations one can check that the bifurcation surface of (E.8) intersects the boundary of the Poincaré coordinates on a sphere of radius $L$ in the boundary CFT with flat background [27]. This sphere is the entangling region for which we want to obtain the bit threads configuration.
The above discussion of coordinate transformations can be extended for a sphere of arbitrary radius $R$ in the CFT [27]. It implies that the horizon (or more specifically the bifurcation surface) of the topological black hole moves to a different position and the intersection with the boundary also changes. In the Poincaré coordinates, the change of position of the bifurcation surface corresponds to the different RT surfaces (arcs) corresponding to the boundary entangling region. One can find a family of arcs which are connected by the isometries of the AdS geometry. A boost in the $(Y_{-1},Y_{2})$ plane of the embedding space is one such isometry. Under this boost, the coordinates transform as
$Y^{\prime}_{-1}=\cosh{\beta}\,Y_{-1}-\sinh{\beta}\,Y_{2},\qquad Y^{\prime}_{2}=\cosh{\beta}\,Y_{2}-\sinh{\beta}\,Y_{-1},$ (E.13)
while keeping the other coordinates remain same. Using these we again end up with the metric (E.8) and choose the Poincaré coordinates to be same as in the (E.6). To relate these two coordinates, we need to write $Y_{-1}$ and $Y_{2}$ in terms of the boost parameter $\beta$ as
$Y_{-1}=\rho\cosh{\beta}\cosh{u}+\tilde{\rho}\sinh{\beta},\qquad Y_{2}=\tilde{\rho}\cosh{\beta}+\rho\sinh{\beta}\cosh{u},$ (E.14)
on a Cauchy surface (i.e., $\tau=0=t$). Now, using the relations (E.5), (E.7) and the above ones, we find
$Y_{-1}+Y_{2}=\dfrac{L^{2}}{z}$ $=\rho\cosh{\beta}\cosh{u}+\tilde{\rho}\sinh{\beta}+\tilde{\rho}\cosh{\beta}+\rho\sinh{\beta}\cosh{u}$
$=\left(\rho\cosh{u}+\tilde{\rho}\right)e^{\beta},$ (E.15)
$Y_{1}=\dfrac{L}{z}x$ $=\rho\sinh{u}\,.$
Similar to the original case, we find the coordinate transformations for the boosted one given by
$\rho=\dfrac{e^{\beta}}{2z}\sqrt{\left(e^{-2\beta}L^{2}+x^{2}+z^{2}\right)^{2}-4e^{-2\beta}L^{2}x^{2}},$ (E.16a)
$u=\operatorname{arctanh}{\left(\dfrac{2e^{-\beta}Lx}{e^{-2\beta}L^{2}+x^{2}+z^{2}}\right)}.$ (E.16b)
The inverse transformations are
$z=\dfrac{e^{-\beta}L^{2}}{\rho\cosh{u}+\tilde{\rho}}\,,\qquad x=\dfrac{e^{-\beta}L\rho\sinh{u}}{\rho\cosh{u}+\tilde{\rho}}$. (E.17)